### v2

# Extremal correlation coefficient for functional data

Abstract
We propose a coefficient that measures extremal dependence in paired samples of functions. It has properties similar to the Pearson correlation, but differs in significant ways: 1) it is designed to measure dependence between curves, 2) it focuses only on extreme curves. The new coefficient is derived within the framework of regular variation in Banach spaces. A consistent estimator is proposed and justified by an asymptotic analysis and a simulation study. The usefulness of the new coefficient is illustrated using financial and climate functional data.
keywords:
Correlation; Extremes; Functional data.
1 Introduction
With the growing impact of extreme events such as financial downturns or unusual weather, there has been increasing interest in developing statistical tools to study patterns of extreme curves. This is to a large extent due to the increasing availability of high resolution data; asset price curves can be constructed at any temporal resolution, and modern weather databases and computer models contain measurements at hourly or even higher frequencies. Such data can be interpreted as curves, e.g., one curve per day, providing more comprehensive view of daily patterns compared to a single summary number like the closing price or maximum temperature. Analyzing extreme curves in the framework of functional data analysis thus leads to a more precise understanding of the impacts associated with extreme events.
This paper makes a methodological and theoretical contribution at the nexus of extreme value theory and functional data analysis. We propose a coefficient that quantifies the tendency of paired extreme curves to exhibit similar patterns simultaneously. Two examples of the type of questions that the tool deals with are the following: 1) During a stock market crisis, such as the market decline due to the COVID-19 pandemic, do returns of different sectors of the economy exhibit similar extreme daily trajectories? 2) How likely is location A to experience a similar daily pattern of temperature as location B (on the same day) during a heat wave? Our proposed coefficient offers a more precise quantification of extreme risk by focusing on 1) the shape of curves and 2) the extreme parts of paired samples. This point is further illustrated with a data example in Section 6.1.
There has been some research focusing on probabilistic and statistical methods for extreme curves. Extreme value theory in the space of continuous functions is studied in Chapters 9 and 10 of de Haan & Ferreira (2000) and Einmahl & Segers (2021). Principal component analysis of extreme curves has been studied by Kokoszka et al. (2019)Kokoszka & Kulik (2023), and Clémençon et al. (2024). Extremal properties of scores of functional data were studied by Kokoszka & Xiong (2018) and Kim and Kokoszka (20192022). Additional, more closely related papers are introduced as we develop our approach. We propose a method for quantifying extremal dependence of paired functional samples, for which there are currently no appropriate tools.
We note that there has been considerable research aimed at quantifying extremal dependence for heavy-tailed random vectors. Ledford and Tawn (1996, 1997, 2003) introduced the coefficient of tail dependence, which was later generalized to the extremogram by Davis & Mikosch (2009). The extremal dependence measure based on the angular measure of a regularly varying random vector was introduced by Resnick (2004) and further investigated by Larsson & Resnick (2012). Janßen et al. (2023) recently introduced a unified approach for representing tail dependence using random exceedence sets. Those measures for extremes are designed for random vectors in a Euclidean space. Therefore, applying any such measures to functional data requires some sort of dimension reduction, e.g., principal component analysis, or data compression like converting daily temperature curves to daily average or maximum values. The reduced data are then analyzed using those tools for multivariate extremes, see, e.g., Meinguet (2010)Dombry & Ribatet (2015), and Kim & Kokoszka (2022). This approach is convenient, but it does not fully utilize all relevant information that functional data contain.
We develop a new measure, the extremal correlation coefficient, that captures the extremal dependence of paired functional samples utilizing the information in the sizes and shapes of the curves. The measure involves an inner product of pairs of extreme curves, and therefore requires finite second moment. Similar ideas have been applied in non-extreme contexts of functional data analysis. Dubin & Müller (2005) introduced a measure, called dynamical correlation, that computes the inner product of all pairs of standardized curves. The concept was further studied by Yeh et al. (2023) where an autocorrelation measure, termed spherical autocorrelation, for functional time series was proposed. These measures are however computed based on the total body of functional data, and so are not suitable for describing extremal dependence.
The coefficient we develop quantifies extremal dependence by specifically focusing on the extreme parts of heavy-tailed functional observations. It is conceptually appealing, as it shares desirable features with the classic correlation coefficient: 1) its values range from -1 to 1, 2) it measures the strength and direction of linear relationship between two extreme curves, 3) if the extremal behavior of two curves is independent, the coefficient is zero. Moreover, it can be used in practice with a relatively simple numerical implementation. We thus hope that such interpretable and tractable tool makes a useful contribution.
Turning to mathematical challenges, the concept of vague convergence, see e.g., Chapters 2 and 6 of Resnick (2007), cannot be readily used. The vague convergence, which now provides a standard mathematical framework for extremes in Euclidean spaces, can be defined only on locally compact spaces. Since every locally compact Banach space has finite dimension, a different framework must be used for functional data in Hilbert spaces. We use the theory of regularly varying measures developed by Hult & Lindskog (2006) who introduced the notion of $M_{0}$ convergence, which works for regularly varying measures on complete separable metric spaces. The $M_{0}$ convergence is further studied by Meinguet (2010), where it is applied to regularly varying time series in a separable Banach space. The concept of $M_{0}$ convergence has been generalized, see Section B.1.1. of Kulik & Soulier (2020), with notable contributions from Lindskog et al. (2014) and Segers et al. (2017). We establish the consistency of the estimator we propose within the framework developed by Hult & Lindskog (2006) and  Meinguet (2010). We proceed through a number of $M_{0}$ convergence results that allow us to apply an abstract Bernstein-type inequality. A method of computing the extremal correlation coefficients analytically (in relatively simple cases) is also developed.
The remainder of the paper is organized as follows. In Section 2, we review regularly varying random elements in Banach spaces. In Section 3, we extend the concept of regular variation to bivariate random elements in a product Banach space. The extremal correlation coefficient is introduced in Section 4, where its asymptotic properties are also studied. Section 5 contains a simulation study, and Section 6 illustrates applications to intraday return curves and daily temperature curves.
Theoretical justification of our approach requires more detailed background and some technical derivations, which are placed in the Supplementary material. Preliminary results for the proof of Theorem 4.4 are presented in Section A, followed by its proof in Section B. In Section C, we provide the proof of Lemma 5.1, and additional numerical results from Sections 5 and 6 are included in Sections D and E, respectively.
2 Regular variation in Banach spaces
[…]
An extreme curve in $L^{2}$ is defined as a functional object with a substantial deviation from the mean function, measured by the $L^{2}$-norm. The norm can be large for various reasons as long as the area under the squares of the curves around the mean function over ${\mathcal{T}}$ is large. For example, curves that are far away from the sample mean or that fluctuate a lot around the sample mean will be extreme according to this definition. Extreme functional observations are thus very different from extreme scalar or multivariate observations because there is a multitude of ways in which a curve can be extreme. We informally call functional data heavy-tailed if the probability that an extreme curve occurs is relatively large.
We now briefly review the $M_{0}$ convergence in a separable Banach space ${\mathbb{B}}$. In what follows, ${0}$ is the zero element. Fix a norm $\|\cdot\|_{{\mathbb{B}}}$ and let $B_{\varepsilon}:=\{z\in{\mathbb{B}}:\|z\|_{{\mathbb{B}}}<\varepsilon\}$ be the open ball of radius $\varepsilon>0$ centered at the origin. A Borel measure $\mu$ defined on ${\mathbb{B}}_{0}:={\mathbb{B}}\setminus\{\mbox{\boldmath${0}$}\}$ is said to be boundedly finite if $\mu(A)<\infty$, for all Borel sets that are bounded away from ${0}$, i.e., $A\cap B_{\varepsilon}=\emptyset$, for some $\varepsilon>0$. Let $M_{0}({\mathbb{B}})$ be the collection of all such measures on ${\mathbb{B}}_{0}$. For $\mu_{n},\mu\in M_{0}({\mathbb{B}})$, the sequence of $\mu_{n}$ converges to $\mu$ in the $M_{0}$ topology ($\mu_{n}\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu$), if $\mu_{n}(A)\to\mu(A)$, for all bounded away from ${0}$, $\mu$–continuity Borel sets $A$, i.e., those with $\mu(\partial A)=0$, where $\partial A$ is the boundary of $A$. Equivalently, $\mu_{n}\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu$, if $\int_{{\mathbb{B}}}f(x)\mu_{n}(dx)\to\int_{{\mathbb{B}}}f(x)\mu(dx)$ for all $f\in{\mathcal{C}}_{0}({\mathbb{B}})$, where ${\mathcal{C}}_{0}({\mathbb{B}})$ is the class of bounded and continuous functions $f:{\mathbb{B}}_{0}\to{\mathbb{R}}$ that vanish on a neighborhood of ${0}$.
We now define regular variation for random elements in ${\mathbb{B}}$, see Theorem 3.1 of Hult & Lindskog (2006) and Chapter 2 of Meinguet (2010). This concept formalizes the idea of heavy-tailed observations in infinite dimensional spaces.
Definition 2.1.
A random element $X$ in ${\mathbb{B}}$ is regularly varying with index $-\alpha$, $\alpha>0$, if there exist a sequence $b(n)\to\infty$ and a measure $\mu$ in $M_{0}({\mathbb{B}})$ such that
$n{\rm pr}\left(\frac{X}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu,\ \ \ \ n\to\infty,$ (1)
where the exponent measure $\mu$ satisfies $\mu(tA)=t^{-\alpha}\mu(A)$ for Borel sets $A\subset{\mathbb{B}}_{0}$.
A possible choice for $b(n)$ is the quantile function, defined by ${\rm pr}(\|X\|_{{\mathbb{B}}}>b(n))=n^{-1}$. Roughly speaking, the tail probability of $X$ decays like a power function, ${\rm pr}(\|X\|_{{\mathbb{B}}}>t)\approx Ct^{-\alpha}$, as $t\to\infty$. The following lemma, see Hult & Lindskog (2006), states an equivalent definition of a regularly varying element in ${\mathbb{B}}$.
Lemma 2.2.
A random element $X$ in ${\mathbb{B}}$ is regularly varying with index $-\alpha$, $\alpha>0$, if and only if there exist a sequence $b^{\prime}(n)\to\infty$ and a probability measure $\Gamma$ on ${\mathbb{S}}:=\{x\in{\mathbb{B}}:\|x\|_{{\mathbb{B}}}=1\}$ (called the angular measure) such that for any $y>0$,
$n{\rm pr}\left(\|X\|_{{\mathbb{B}}}>b^{\prime}(n)y,X/\|X\|_{{\mathbb{B}}}\in\cdot\right)\stackrel{{\scriptstyle w}}{{\longrightarrow}}cy^{-\alpha}\Gamma,\ \ \ \ n\to\infty,$ (2)
for some $c>0$.
[…]
Figure 1: The first three orthonormal basis elements in $L^{2}[0,1]$ defined in (3) (left-most); simulated data when $\Gamma$ concentrates on $\phi_{1}$(second from the left); on $\phi_{2}$ (third left); on $\phi_{3}$ (fourth left).
If Definition 2.1 (or condition (2)) holds, we write $X\in RV(-\alpha,\Gamma)$. The polar representation (2) provides an intuitive interpretation of regular variation in ${\mathbb{B}}$. It characterizes regular variation of $X$ in ${\mathbb{B}}$ using two components, the tail index $\alpha$ and the angular probability measure $\Gamma$. The tail index $\alpha$ quantifies how heavy the tail distribution of $\|X\|_{{\mathbb{B}}}$ is, e.g., the probability of extreme curves occurring gets higher as $\alpha$ gets smaller. While the tail index $\alpha$ determines the frequency of occurrence of extreme curves, the angular measure $\Gamma$, defined on the unit sphere ${\mathbb{S}}$, fully characterizes the distribution of the shape of the scaled extreme curves, $X/\|X\|_{{\mathbb{B}}}$. To illustrate this, consider a set of orthonormal functions in $L^{2}([0,1])$ of the form
$\phi_{j}(t)=\sqrt{2}\sin\left(\left(j-\frac{1}{2}\right)\pi t\right),\ \ \ \ j=1,2,\ldots,\ \ t\in[0,1].$ (3)
The first three functions are shown in the left-most plot of Figure 1. We consider a finite-dimensional subspace of $L^{2}([0,1])$, spanned by the first 9 $\phi_{j}$’s, for the purpose of simulations. The data generating process is $X(t)=\sum_{j=1}^{9}Z_{j}\phi_{j}(t)$, where $[Z_{1},\ldots,Z_{9}]^{\top}$ is a 9-dimensional random vector with independent components. Suppose that $Z$ is a random variable following a Pareto distribution with tail index $\alpha=3$ and $N$ is a normal random variable with mean 0 and variance 0.5. We consider the following three cases for $[Z_{1},\ldots,Z_{9}]^{\top}$:
1. $[Z,N,N,N,\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{1}$.
2. $[N,Z,N,N\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{2}$.
3. $[N,N,Z,N\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{3}$.
In all three cases, it follows from Proposition 7.1 and Example 7.3 of Meinguet & Segers (2010) that $X(t)$ is regularly varying with tail index $\alpha=3$. Figure 1 displays simulated data with sample size of 100 for each of the three cases. The plots of simulated data clearly show that the angular measure $\Gamma$ represents the distribution of the shapes of extreme curves in that they are dominated by the shape of the functional axis $\phi_{j}$ on which $\Gamma$ concentrates.
3 Bivariate regular variation in Banach spaces
In order to describe the extremal dependence of two regularly varying random elements $X$ and $Y$ in $L^{2}$, we need to identify their joint probabilistic behavior. We again study it in the more general space ${\mathbb{B}}^{2}$. We propose the following definition.
Definition 3.1.
A bivariate random element $[X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ is said to be jointly regularly varying with index $-\alpha$, $\alpha>0$, if there exist a sequence $b(n)\to\infty$ and a measure $\mu$ in $M_{0}({\mathbb{B}}^{2})$ such that
$n{\rm pr}\left(\frac{(X,Y)}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu,\ \ \ \ n\to\infty,$ (4)
where the joint exponent measure $\mu$ satisfies $\mu(tA)=t^{-\alpha}\mu(A)$ for Borel sets $A\subset{\mathbb{B}}_{0}^{2}$.
We assume that one-dimensional marginal distributions of $\mu$ are non-degenerate, i.e., $\mu_{X}:=\mu(\cdot\times{\mathbb{B}})$ and $\mu_{Y}:=\mu({\mathbb{B}}\times\cdot)$ are measures in $M_{0}({\mathbb{B}})$ satisfying analogs of (1). Since $X$ and $Y$ are normalized by the same function $b(n)$, the marginal distributions are tail equivalent. A possible choice for $b(n)$ is the quantile function, defined by
${\rm pr}(\|(X,Y)\|_{{\mathbb{B}}^{2}}>b(n))=n^{-1}.$
With this choice, we have that
$n{\rm pr}\left(\frac{(X,Y)}{b(n)}\in{\mathcal{A}}_{1}\right)=\frac{{\rm pr}\left((X,Y)\in b(n){\mathcal{A}}_{1}\right)}{{\rm pr}(\|(X,Y)\|_{{\mathbb{B}}^{2}}>b(n))}=\frac{{\rm pr}\left((X,Y)\in{\mathcal{A}}_{b(n)}\right)}{{\rm pr}\left((X,Y)\in{\mathcal{A}}_{b(n)}\right)}=1,$
where ${\mathcal{A}}_{r}$ is defined by
$\mathcal{A}_{r}=\{(x,y)\in{\mathbb{B}}^{2}:\|(X,Y)\|_{{\mathbb{B}}^{2}}\geq r\},\ \ \ \ r>0.$ (5)
Thus, it follows from the $M_{0}$ convergence in (4) and Lemma A.1 that
$\mu({\mathcal{A}}_{1})=\mu\{(x,y)\in{\mathbb{B}}^{2}:\|(X,Y)\|_{{\mathbb{B}}^{2}}>1\}=1,$ (6)
which implies that $\mu$ is a probability measure on ${\mathcal{A}}_{1}$. Throughout the paper, we set
$\|(x,y)\|_{{\mathbb{B}}^{2}}:=\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}.$
This choice of norm works well with the extremal correlation coefficient defined in Section 4.
In order to derive the joint angular probability measure of $X$ and $Y$, we consider the polar coordinate transformation $T:{\mathbb{B}}_{0}^{2}\to\left([0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}\right)\times{\mathbb{S}}^{2}$, defined by
$T(x,y)=\left(\|x\|_{{\mathbb{B}}},\|y\|_{{\mathbb{B}}},\frac{x}{\|x\|_{{\mathbb{B}}}},\frac{y}{\|y\|_{{\mathbb{B}}}}\right)=:(r_{X},r_{Y},\theta_{X},\theta_{Y}),\ \ \ \ (x,y)\in{\mathbb{B}}_{0}^{2}.$ (7)
Using $T$, we obtain an equivalent formulation for a regularly varying random element in ${\mathbb{B}}^{2}$.
Proposition 3.2.
A bivariate random element $[X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ is regularly varying with index $-\alpha$, $\alpha>0$, if and only if there exists an exponent measure $\nu$ in $M_{0}([0,\infty)^{2})$ and an angular probability measure $\Gamma$ in $M_{0}({\mathbb{S}}^{2})$ such that
$n{\rm pr}\left(\frac{(\|X\|_{{\mathbb{B}}},\|Y\|_{{\mathbb{B}}})}{b(n)}\in\cdot,\left(X/\|X\|_{{\mathbb{B}}},Y/\|Y\|_{{\mathbb{B}}}\right)\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\nu\times\Gamma,\ \ n\to\infty,$ (8)
where $b(n)$ is the increasing sequence in (4). (We note that $\mu$ in (4) is a measure on ${\mathbb{B}}_{0}^{2}$, while $\nu$ in (8) is a measure on $[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}$.)
Proof 3.3.
We only show that (4) implies (8) since showing the converse is similar. Take any $f\in{\mathcal{C}}_{0}([0,\infty)^{2}\times{\mathbb{S}}^{2})$. It then follows from the change of variables that
$\int_{[0,\infty)^{2}}\int_{{\mathbb{S}}^{2}}f(r_{X},r_{Y},\theta_{X},\theta_{Y})\ n{\rm pr}\left(\frac{(\|X\|_{{\mathbb{B}}},\|Y\|_{{\mathbb{B}}})}{b(n)}\in(dr_{X},dr_{Y}),\left(\frac{X}{\|X\|_{{\mathbb{B}}}},\frac{Y}{\|Y\|_{{\mathbb{B}}}}\right)\in(d\theta_{X},d\theta_{Y})\right)$
$=\int_{{\mathbb{B}}^{2}}f(T(x,y))\ n{\rm pr}\left(\frac{X}{b(n)}\in dx,\frac{Y}{b(n)}\in dy\right).$
Since $f\in{\mathcal{C}}_{0}([0,\infty)^{2}\times{\mathbb{S}}^{2})$, there exists a set $A$, bounded away from ${0}$ in $[0,\infty)^{2}\times{\mathbb{S}}^{2}$, such that $f(T(x,y))=0$ if $T(x,y)\notin A$. Then we have that $f(T(x,y))=0$ if $(x,y)\notin T^{-1}(A)$. Since $T^{-1}(A)$ is bounded away from ${0}$ in ${\mathbb{B}}^{2}$, we have that $f\circ T\in{\mathcal{C}}_{0}({\mathbb{B}}^{2})$. Then by the $M_{0}$ convergence in Definition 3.1, we have that
$\int_{{\mathbb{B}}^{2}}(f\circ T)(x,y)\ n{\rm pr}\left(\frac{X}{b(n)}\in dx,\frac{Y}{b(n)}\in dy\right)$
$\to\int_{{\mathbb{B}}^{2}}(f\circ T)(x,y)\mu(dx,dy)=\int_{T({\mathbb{B}}^{2})}f(r_{X},r_{Y},\theta_{X},\theta_{Y})\mu\circ T^{-1}(dr_{X},dr_{Y},d\theta_{X},d\theta_{Y}).$
To investigate the form of $\mu\circ T^{-1}$, take any $t>0$ and Borel set $S\subset{\mathbb{S}}^{2}$. It then follows from the homogeneity property of $\mu$ that
$\mu\circ T^{-1}([0,\infty)^{2}\setminus[0,t]^{2}\times S)$
$=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>t,(x/\|x\|_{{\mathbb{B}}},y/\|y\|_{{\mathbb{B}}})\in S\right\}$
$=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>t\right\}\times$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{t^{-\alpha}\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>1,(x/\|x\|_{{\mathbb{B}}},y/\|y\|_{{\mathbb{B}}})\in S\right\}}{t^{-\alpha}\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>1\right\}}.$
It then follows from (6) that
$\mu\circ T^{-1}([0,\infty)^{2}\setminus[0,t]^{2}\times S)=\nu([0,\infty)^{2}\setminus[0,t]^{2})\Gamma(S),$
where
$\nu(A):=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:(\|x\|_{{\mathbb{B}}},\|y\|_{{\mathbb{B}}})\in A\right\},\ \ \ \ A\subset[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\};$
$\Gamma(S):=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>1,(x/\|x\|_{{\mathbb{B}}},y/\|y\|_{{\mathbb{B}}})\in S\right\},\ \ \ \ S\subset{\mathbb{S}}^{2}.$
Thus, $\mu\circ T^{-1}$ has the product form such that on $([0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\})\times{\mathbb{S}}^{2}$
$\mu\circ T^{-1}=\nu\times\Gamma,$ (9)
which completes the proof.
Convergence (8) enables characterizing the joint extremal behavior of the bivariate regularly varying random vector $[X,Y]^{\top}$ in ${\mathbb{B}}^{2}$, using two measures: $\nu$ on $[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}$ and $\Gamma$ on ${\mathbb{S}}^{2}$. The measure $\nu$ describes the joint extremal behavior of the scalars $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$. Specifically, if $\nu$ has its mass only on the axes, then $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$ are asymptotically independent, i.e., if one curve shows an extreme pattern, there is negligible probability of the other curve also showing an extreme pattern. If $\nu$ has mass only on the line $\{t(1,1),t>0\}$, then $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$ show asymptotic full dependence, i.e., extreme curves occur simultaneously in $X$ and $Y$.
The difference between $\nu$ in (8) and $\mu$ in (4) is that $\mu$ describes the joint behavior of extreme curves $X$ and $Y$ in ${\mathbb{B}}_{0}^{2}$, but $\nu$ describes extremal dependence between the sizes $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$ in $[0,\infty)^{2}\setminus\mbox{\boldmath${0}$}$. We remark that the measure $\nu$ on $[0,\infty)^{2}\setminus\mbox{\boldmath${0}$}$ is homogeneous like $\mu$ because for any $A\subset[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}$ and $t>0$,
$\nu(tA)=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:(\|x\|_{{\mathbb{B}}},\|y\|_{{\mathbb{B}}})\in tA\right\}=t^{-\alpha}\nu(A).$
The joint angular probability measure $\Gamma$ characterizes how the shapes of scaled $X$ and $Y$ are related in extremes. If the extreme curves are exactly proportional, i.e., $X=\lambda Y$, $\lambda\neq 0$, the scaled curves share the same extreme functional elements. This means that $\Gamma$ concentrates on the “line” $\{(\phi_{j},\phi_{j}),\ j\in{\mathcal{J}}\subset{\mathbb{N}}\}\subset{\mathbb{S}}^{2}$, where $\{\phi_{j},\ j\geq 1\}$ is a set of orthonormal elements in ${\mathbb{S}}$. If the shapes of two curves do not match in extremes, then $\Gamma$ concentrates on $\{(\phi_{j},\phi_{k})\}\subset{\mathbb{S}}^{2}$, where $j\in{\mathcal{J}}\subset{\mathbb{N}}$, $k\in{\mathcal{K}}\subset{\mathbb{N}}$, ${\mathcal{J}}\cap{\mathcal{K}}=\emptyset$. This situation corresponds to vanishing extremal covariance defined in Section 4.
The marginal extreme behavior of $X$ can be obtained by integrating all possible values of $Y$ in (4), or $\|Y\|_{{\mathbb{B}}},Y/\|Y\|_{{\mathbb{B}}}$ in (8). Then $X$ has its marginal measure $\mu_{X}$, and equivalently $X\in RV(-\alpha,\Gamma_{X})$, where $\Gamma_{X}$ is the marginal angular measure of $X$. Similarly, $Y$ has its marginal $\mu_{Y}$, and equivalently $Y\in RV(-\alpha,\Gamma_{Y})$, where $\Gamma_{Y}$ is the marginal angular measure of $Y$.
4 Extremal correlation coefficient for functional data
In this section, we introduce the extremal correlation coefficient for functional data. It focuses on the extreme part of the joint distribution of regularly varying random elements $X$ and $Y$ in $L^{2}$ and measures the tendency of paired curves to exhibit similar extreme patterns.
To define the extremal correlation coefficient, we begin by introducing the concept of extremal covariance. Given a regularly varying bivariate random element $[X,Y]^{\top}$ in $L^{2}\times L^{2}$ with joint exponent measure $\mu$ in (4), we define the extremal covariance between $X$ and $Y$ by
$\sigma_{XY}=\int_{\|x\|\vee\|y\|>1}\left\langle x,y\right\rangle\mu(dx,dy).$ (10)
Recall that by (6), $\mu$ is a probability measure on the domain $\{(x,y)\in L^{2}\times L^{2}:\|x\|\vee\|y\|>1\}$, so $\mu I_{\|x\|\vee\|y\|>1}$ represents a probability distribution describing the joint extremal behavior of $X$ and $Y$. The extremal covariance is thus an extreme analog of the classic covariance in that $\sigma_{XY}$ measures how much two random curves vary together in extremes. We note that in order to define extremal covariance in (10), $[X,Y]^{\top}$ must be regularly varying with index $-\alpha$, where $\alpha>2$. The condition $\alpha>2$ is necessary because the definition of extremal covariance presumes the existence of the second moment, just as the usual covariance does. We elaborate on it at the end of this section, but note here that, as explained at the beginning of Section 5, regularly varying functions can be transformed to have the index $\alpha>2$.
Table 1: The range of $\sigma_{XY}$, extremal covariance between $X$ and $Y$, depending on the extremal dependence between $\|X\|$ and $\|Y\|$ and the level of similarity between $X/\|X\|$ and $Y/\|Y\|$ in extremes.
$\|X\|$ and $\|Y\|$ $\|X\|$ and $\|Y\|$
Asymptotic Independence Asymptotic Dependence
$X/\|X\|$ and $Y/\|Y\|$
$\sigma_{XY}=0$     look similar $\sigma_{XY}>0$
    look orthogonal $\sigma_{XY}\approx 0$
    look opposite $\sigma_{XY}<0$
To examine the extremal covariance more closely, we recall the transformation $T$ defined in (7) and the relation $\mu\circ T^{-1}=\nu\times\Gamma$ in (9). It then follows from the change of variables that
$\sigma_{XY}=\int_{r_{X}\vee r_{Y}>1}r_{X}r_{Y}\ \nu(dr_{X},dr_{Y})\int_{{\mathbb{S}}^{2}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y}).$ (11)
The extremal covariance of $X$ and $Y$ can be thus factorized into two components: the extremal dependence between $\|X\|$ and $\|Y\|$ (the first factor in (11)) and the level of similarity of the shapes between $X/\|X\|$ and $Y/\|Y\|$ (the second factor in (11)). These two components can be considered separately, each of them contains useful information, but there is value in combining them. If $\|X\|$ and $\|Y\|$ are asymptotically (in extremes) independent, then $\sigma_{XY}=0$ regardless of how the (scaled) extreme shapes of $X$ and $Y$ compare. The coefficient $\sigma_{XY}$ thus preserves the property that bivariate scalar observations are independent in extremes if they cannot take extremely large values at the same time. If “large” curves tend to occur simultaneously, i.e., $\|X\|$ and $\|Y\|$ are asymptotically dependent, then $\int_{r_{X}\vee r_{Y}>1}r_{X}r_{Y}\ \nu(dr_{X},dr_{Y})>0$. Consequently, there are three possible ranges for $\sigma_{XY}$, depending on the relative shape of extreme $X/\|X\|$ and $Y/\|Y\|$: 1) $\sigma_{XY}>0$ if the shapes look similar, 2) $\sigma_{XY}\approx 0$ if the shapes do not match, i.e., the curves are orthogonal, or 3) $\sigma_{XY}<0$ if they look opposite. These properties are summarized in Table 1. To provide a scale invariant measure, we define the extremal correlation coefficient.
Definition 4.1.
The extremal correlation coefficient is defined by
$\rho_{XY}=\frac{\sigma_{XY}}{\sigma_{X}\sigma_{Y}},$ (12)
where the extremal covariance $\sigma_{XY}$ is defined by (10), and where
$\sigma_{X}=\left\{\int_{\|x\|\vee\|y\|>1}\|x\|^{2}\mu(dx,dy)\right\}^{1/2},\ \ \ \ \sigma_{Y}=\left\{\int_{\|x\|\vee\|y\|>1}\|y\|^{2}\mu(dx,dy)\right\}^{1/2}.$
The coefficient $\rho_{XY}$ has properties analogous to the classic correlation coefficient: 1) $-1\leq\rho_{XY}\leq 1$, 2) $\rho_{XY}$ measures the strength and direction of linear relationships between $X$ and $Y$ in extremes, 3) If $X$ and $Y$ are independent, then $\rho_{XY}=0$ since independence implies asymptotic independence between $\|X\|$ and $\|Y\|$.
To motivate our estimation approach, we first show that $\sigma_{XY}$ is a limit of the expected inner product of $X$ and $Y$ conditional on large values of $[X,Y]^{\top}$.
Proposition 4.2.
Let $[X,Y]^{\top}$ be a regularly varying random element in $L^{2}\times L^{2}$. Then,
$\sigma_{XY}=\lim_{n\to\infty}E\left[\left\langle\frac{X}{b(n)},\frac{Y}{b(n)}\right\rangle\Bigg{|}\|X\|\vee\|Y\|>b(n)\right].$
Proof 4.3.
Considering $f:L^{2}\times L^{2}\to{\mathbb{R}}$ defined by $(x,y)\to\left\langle x,y\right\rangle I_{\|x\|\vee\|y\|>1}$, we have that
$E\left[\left\langle{b(n)}^{-1}X,{b(n)}^{-1}Y\right\rangle|\|X\|\vee\|Y\|>{b(n)}\right]$
$=\frac{1}{{\rm pr}(\|X\|\vee\|Y\|>{b(n)})}E\left[\left\langle{b(n)}^{-1}X,{b(n)}^{-1}Y\right\rangle I_{\|X\|\vee\|Y\|>{b(n)}}\right]$
$=\int_{L^{2}\times L^{2}}f(x,y)\frac{{\rm pr}({b(n)}^{-1}X\in dx,{b(n)}^{-1}Y\in dy)}{{\rm pr}(\|X\|\vee\|Y\|>{b(n)})}.$
Note that $f$ is bounded and vanishes on a neighborhood of ${0}$ in $L^{2}\times L^{2}$. Also, the discontinuity set of $f$ is the boundary of $\mathcal{A}_{1}=\{(x,y)\in L^{2}\times L^{2}:\|x\|\vee\|y\|\geq 1\}$, and it follows from Lemma A.1 that $\mu(\partial\mathcal{A}_{1})=0$. Therefore, by (4) and Lemma A.1 of Meinguet & Segers (2010), we get the claim.
Based on Proposition 4.2, we propose an estimator for $\sigma_{XY}$ defined by
$\hat{\sigma}_{n,k}=\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{R_{(k)}},\frac{Y_{i}}{R_{(k)}}\right\rangle I_{R_{i}\geq R_{(k)}},$ (13)
where $[X_{i},Y_{i}]^{\top},1\leq i\leq n$, are i.i.d. copies of $[X,Y]^{\top}$, $R_{i}:=\|X_{i}\|\vee\|Y_{i}\|$ and $R_{(k)}$ is the $k$th largest order statistic with the convention $R_{(1)}=\max\{R_{1},\ldots,R_{n}\}$. An estimator for $\rho_{XY}$ is then defined by
$\hat{\rho}_{n,k}=\frac{\sum_{i=1}^{n}\left\langle X_{i},Y_{i}\right\rangle}{\sqrt{\sum_{i=1}^{n}\|X_{i}\|^{2}}\sqrt{\sum_{i=1}^{n}\|Y_{i}\|^{2}}}\ I_{R_{i}\geq R_{(k)}}.$ (14)
These estimators take only the $k$ largest pairs of $[X_{i},Y_{i}]^{\top}$, $1\leq i\leq n$, based on their norm, i.e., $\|X_{i}\|\vee\|Y_{i}\|$, as inputs. This approach falls into so-called peaks-over-threshold framework in that it relies only on $k$ largest observations whose magnitude exceeds a certain threshold. Asymptotic properties in this framework are typically derived as $k$ goes to infinity with $n$, in such a way that $k/n\to 0$. We assume throughout the paper that this condition holds.
We will work under the following assumption.
{assumption}
The bivariate random element $[X,Y]^{\top}$ in $L^{2}\times L^{2}$ has mean zero and is regularly varying with index $-\alpha$, $\alpha>2$. The observations $[X_{1},Y_{1}]^{\top},[X_{2},Y_{2}]^{\top},\ldots$ are independent copies of $[X,Y]^{\top}$.
We state in the following theorem that the estimator $\hat{\sigma}_{n,k}$ is consistent for the extremal covariance. All proofs of the theoretical results introduced in this section are presented in Sections A and B of the Supplementary material, as they require a number of preliminary results and technical arguments.
Theorem 4.4.
Under Assumption 4, $\hat{\sigma}_{n,k}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\sigma_{XY}$, where $\hat{\sigma}_{n,k}$ and $\sigma_{XY}$ are defined in (13) and (10), respectively.
From Theorem 4.4, the consistency of $\hat{\rho}_{n,k}$ for $\rho_{XY}$ follows from Slutsky’s theorem.
Corollary 4.5.
Under Assumption 4, $\hat{\rho}_{n,k}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\rho_{XY}$, where $\hat{\rho}_{n,k}$ and $\rho_{XY}$ are defined in (14) and (12), respectively.
We end this section with a discussion on the condition $\alpha>2$ in Assumption 4, which ensures the existence of the second moments, $E\|X\|^{2}$ and $E\|Y\|^{2}$. It can be dropped for the following alternative measure
$\gamma_{XY}:=\int_{{\mathbb{S}}^{2}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y}),$
which is the angular density factor in (11). An estimator for $\gamma_{XY}$ can be defined by $\hat{\gamma}_{n,k}=\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{\|X_{i}\|},\frac{Y_{i}}{\|Y_{i}\|}\right\rangle I_{R_{i}\geq R_{(k)}}$, and its consistency can be proven in almost the same manner as the proof of Corollary 4.2 of Clémençon et al. (2024). While $\gamma_{XY}$ can measure the similarity of curve shapes, it does not account for whether extreme curves $X$ and $Y$ occur simultaneously. Therefore, it can be argued that $\rho_{XY}$ is a more appropriate measure for evaluating concurrent risk despite requiring $\alpha>2$, as it assesses the similarity of shapes in paired extreme curves that occur simultaneously. To see this more closely, suppose that $X=Z_{1}\phi$ and $Y=Z_{2}\phi$, where $Z_{1}$, $Z_{2}$ are i.i.d. random variables satisfying ${\rm pr}(Z_{1}>z)=z^{-\alpha}$, $z>0$, and $\phi$ is any unit norm element of $L^{2}$. In this setting, two curves $X$ and $Y$ will appear identical in their extremes, but the extreme curves are unlikely to be observed simultaneously due to the independence of $Z_{1}$ and $Z_{2}$. Therefore, the assessment of concurrent risk should be zero, and the measure that gives zero is $\rho_{XY}=0$ (while $\gamma_{XY}=1)$. The identity $\rho_{XY}=0$ holds because the measure $\nu$ in (11) concentrates on the coordinate axes, so $\sigma_{XY}=\int_{r_{X}\vee r_{Y}>1}r_{X}r_{Y}\ \nu(dr_{X},dr_{Y})\int_{{\mathbb{S}}^{2}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y})=0\times 1=0$, and the condition $\alpha>2$ is needed for the first integral to exist.
5 A simulation study
In this simulation study, we demonstrate that the proposed estimator $\hat{\rho}_{n,k}$ consistently estimates the extremal correlation coefficient. Before proceeding, we explain how $\hat{\rho}_{n,k}$ can be computed when curves are observed at discrete points. This will be applied throughout Sections 5 and 6.
Assume that curves are observed on the regularly spaced grid $\{j/J,\ j\in\{1,\ldots,J\}\}$ on $[0,1]$, with each point assigned equal weight. The inner product and norm in $L^{2}([0,1])$ are then given by
$\left\langle x,y\right\rangle=\frac{1}{J}\sum_{j=1}^{J}x(j/J)y(j/J),\quad\|x\|=\left[\frac{1}{J}\sum_{j=1}^{J}x(j/J)^{2}\right]^{1/2},\quad x,y\in L^{2}([0,1]).$ (15)
Even if the curves are observed at irregularly spaced or different grids, or contain missing values, they can be reconstructed on a regularly spaced grid, see e.g., Chapters 1 and 7 in Kokoszka & Reimherr (2017). Using (15), $\hat{\rho}_{n,k}$ can be computed in the following steps:
Step 1. Verify if $\|X\|$ and $\|Y\|$ are regularly varying, for example, by examining whether their Hill plots exhibit stable regions.
Step 2. Estimate the tail indexes of $\|X\|$ and $\|Y\|$ using the Hill estimator. For this, we use the mindist function from the R package tea.
Step 3. If the tail index estimates from Step 2 are not close to each other, apply a transformation to make $X$ and $Y$ tail equivalent. One approach is the power transformation discussed on page 310 of Resnick (2007). Given $X\in RV(-\alpha_{X},\Gamma_{X})$ and $Y\in RV(-\alpha_{Y},\Gamma_{Y})$, consider the transformation
$g_{X}(x)=\frac{x}{\|x\|^{1-{\alpha_{X}}/\alpha}},\ \ \ \ g_{Y}(y)=\frac{y}{\|y\|^{1-{\alpha_{Y}}/\alpha}},\ \ \ \ x,y\in L^{2},$ (16)
where $\alpha$ is a desired tail index. Applying $g_{X}$ and $g_{Y}$ to $X$ and $Y$, respectively, ensures that $P(\|g_{X}(X)\|>\cdot)$ and $P(\|g_{Y}(Y)\|>\cdot)$ are regularly varying with the same index $-\alpha$. Since this method adjusts only the scale of curves, the transformed curves still retain their original shapes.
Step 4. Given tail equivalent marginals, take the $k$ largest pairs of $[X_{i},Y_{i}]^{\top}$, based on their norm, i.e., $R_{i}=\|X_{i}\|\vee\|Y_{i}\|$, and then compute $\hat{\rho}_{n,k}$ as in (14). By Lemma A.3 (i), if $[X,Y]^{\top}$ are regularly varying in $L^{2}\times L^{2}$ with index $-\alpha$, then $R=\|X\|\vee\|Y\|$ is regularly varying in $\mathbb{R}_{+}$ with the same index. Therefore, we choose $k$ that results in successful tail estimation for $R$ in finite samples. In the literature on tail estimation, various methods for selecting $k$ have been introduced, e.g., Hall & Welsh (1985)Hall (1990)Drees & Kaufmann (1998),  Danielsson et al. (2001), just to name a few. We use the method of  Danielsson et al. (2016), which is implemented using the function mindist of the R package tea.
We now outline the design of our simulation study. We generate functional observations in such a way that the theoretical value of $\rho_{XY}$ can be computed analytically, so that we can see how close $\hat{\rho}_{n,k}$ is to the true value. Suppose that $Z_{1}$ and $Z_{2}$ are i.i.d. random variables in ${\mathbb{R}}$ satisfying ${\rm pr}(|Z_{1}|>z)=z^{-\alpha}$ with equal chance of $Z_{1}$ being either negative or positive. Also, let $N_{1}$, $N_{2}$, and $N_{3}$ be i.i.d. normal random variables in ${\mathbb{R}}$ with mean 0 and variance 0.5. Consider $\{\phi_{j},j\geq 1\}$ defined by (3) and recall that it is an orthonormal basis in $L^{2}([0,1])$. These functions are simulated on a grid of 100 equally–spaced points on the unit interval $[0,1]$. We consider the following data generating processes, for $-1\leq\rho\leq 1$,
$X(t)=Z_{1}\phi_{1}(t)+N_{1}\phi_{2}(t)+N_{2}\phi_{3}(t);\ \ Y(t)=\rho Z_{1}\phi_{1}(t)+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}(t)+N_{3}\phi_{3}(t).$ (17)
It generates extreme curves dominated by the shape of the functional axis $\phi_{1}$ for $X$ and by either $\phi_{1}$ or $\phi_{2}$ for $Y$. The following lemma gives an analytic formula for $\rho_{XY}$. Its proof (and a slightly more general result) is provided in Section C of Supplementary material.
Lemma 5.1.
Let $[Z_{1},Z_{2}]^{\top}$ be a random vector in ${\mathbb{R}}^{2}$ consisting of i.i.d. components $Z_{j}$ whose magnitude is regularly varying with tail index $\alpha>2$, i.e., for some $c_{+}$, $c_{-}\geq 0$, ${\rm pr}(Z_{1}>z)\sim c_{+}z^{-\alpha}$ and ${\rm pr}(Z_{1}<-z)\sim c_{-}z^{-\alpha}$, where $f(z)\sim g(z)$ if and only if $\lim_{z\to\infty}f(z)/g(z)=1$. Also, let $\{\phi_{j},\ j\geq 1\}$ be a set of orthonormal elements in ${\mathbb{S}}$. Then, for $X$ and $Y$ in (17),
$\rho_{XY}=\frac{\rho}{\{\rho^{2}+(1-\rho^{2})^{\alpha/2}\}^{1/2}}.$
We consider $\rho_{XY}\in\{0,\pm 0.1,\pm 0.2,\ldots,\pm 0.9,\pm 1\}$ and $\alpha\in\{3,4,5\}$, from which values of $\rho$ can be obtained by Lemma 5.1. For each $\rho$, we generate $[X_{i},Y_{i}]^{\top}$, $1\leq i\leq n$, that are i.i.d. copies of $[X,Y]^{\top}$, with sample sizes $n\in\{100,500,2000\}$. In each case, 1000 replications are generated. By Proposition 7.1 and Example 7.3 of Meinguet & Segers (2010), $X$ and $Y$ in (17) are regularly varying with the same index $-\alpha$, so we proceed directly to Step 4 and compute $\hat{\rho}_{n,k}$. When choosing $k$, we also consider an alternative approach based on the KS distance, as introduced by Clauset et al. (2009). The key difference from Danielsson et al. (2016) is that Clauset et al. (2009) computes the distance from tail distributions, rather than tail quantiles. The method is implemented using poweRlaw R package.
Table 2: The magnitude of empirical biases (standard errors) of $\hat{\rho}_{n,k}$ when $\alpha=3$. Optimal $k$s are selected using the method from Danielsson et al. (2016), with averages of $k=8$ ($N=100$), $k=26$ ($N=500$), and $k=63$ ($N=2000$).
$\rho_{XY}$ $N=100$ $N=500$ $N=2000$
-1.0 0.04 (0.03) 0.02 (0.01) 0.01 (0.01)
-0.9 0.06 (0.13) 0.04 (0.09) 0.03 (0.09)
-0.8 0.07 (0.16) 0.04 (0.13) 0.03 (0.12)
-0.7 0.04 (0.18) 0.02 (0.15) 0.01 (0.14)
-0.6 0.03 (0.19) 0.02 (0.16) 0.01 (0.14)
-0.5 0.01 (0.19) 0.00 (0.16) 0.01 (0.15)
-0.4 0.01 (0.18) 0.00 (0.15) 0.01 (0.15)
-0.3 0.01 (0.16) 0.01 (0.13) 0.02 (0.14)
-0.2 0.02 (0.14) 0.02 (0.12) 0.01 (0.11)
-0.1 0.01 (0.09) 0.01 (0.08) 0.01 (0.09)
0.0 0.00 (0.06) 0.00 (0.04) 0.00 (0.03)
0.1 0.01 (0.10) 0.00 (0.08) 0.01 (0.09)
0.2 0.01 (0.14) 0.02 (0.11) 0.02 (0.11)
0.3 0.02 (0.16) 0.02 (0.15) 0.01 (0.13)
0.4 0.00 (0.18) 0.02 (0.15) 0.01 (0.15)
0.5 0.00 (0.19) 0.00 (0.16) 0.00 (0.15)
0.6 0.04 (0.18) 0.02 (0.16) 0.00 (0.15)
0.7 0.04 (0.17) 0.03 (0.15) 0.01 (0.14)
0.8 0.06 (0.16) 0.03 (0.12) 0.03 (0.13)
0.9 0.06 (0.12) 0.04 (0.10) 0.03 (0.09)
1.0 0.04 (0.03) 0.02 (0.01) 0.01 (0.01)
We report the magnitude of empirical biases (the absolute difference between the average and the theoretical value), along with standard errors computed as the sample standard deviations. Using the optimal $k$s selected by the method from Danielsson et al. (2016), the results are shown in Table 2 for $\alpha=3$. The results for $\alpha\in\{4,5\}$, as well as those using the method from Clauset et al. (2009), are provided in Section D.1 of Supplementary material, as they lead to similar conclusions. The conclusions are summarized as follows.
1. The estimator is consistent as the biases decrease with increasing sample sizes, across nearly all values of $\rho_{XY}$ and for all $\alpha\in\{3,4,5\}$ considered. However, the finite performance using the optimal $k$s appears to depend on the tail index $\alpha$, with smaller $\alpha$ resulting in small biases.
2. The bias tends to increase in magnitude as $|\rho_{XY}|$ approaches 1. This could be due to the effect of the boundary, $\rho_{XY}\in\{-1,1\}$. These barriers cause the bias to increase, although it gets lower again at the boundary.
3. The standard errors are observed to be non-uniform across $\rho_{XY}$, they roughly behave like a quadratic function of $\rho_{XY}$ with its peak at ±.$\pm.5$.
Additional information about $\hat{\rho}_{n,k}$ is provided in the online Supplementary material. Specifically, Section D.1 examines the effect of the selection of $k$, Section D.2 explores the relation to the extremal measures $\chi$ and $\bar{\chi}$ introduced by Coles et al. (1999), and Section D.3 investigates the effect of phase variation.
6 Applications to financial and climate functional data
6.1 Extremal dependence of intraday returns on sector ETFs
In this section, we study pairwise extremal dependence of cumulative intraday return curves (CIDRs) of Exchange Traded Funds (ETFs) reflecting performance of key sectors of the U.S. economy. We work with nine Standard & Poor’s Depositary Receipt ETFs listed in Table 3. Our objective is to measure the tendency of paired CIDRs to exhibit similar extreme daily trajectories during the market decline caused by the COVID-19 pandemic. The CIDRs are defined as follows. Denote by $P_{i}(t)$ the price of an asset on trading day $i$ at time $t$. For the assets in our example, $t$ is time in minutes between 9:30 and 16:00 EST (NYSE opening times) rescaled to the unit interval $(0,1)$. The CIDR on day $i$ is the curve $R_{i}(t)=\ln P_{i}(t)-\ln P_{i}(0)$, $t\in[0,1]$, where $P_{i}(0)$ is the opening price on day $i$. The curves $R_{i}$ show how the return accumulates over the trading day, see Figure 2. We consider all full trading days between Jan 02, 2020 and July 31, 2020 ($N=147$).
Table 3: The nine sector ETFs and their corresponding tail index estimates $\hat{\alpha}$.
Ticker Sector $\hat{\alpha}$
XLY Consumer Discretionary 3.8
XLP Consumer Staples 2.6
XLE Energy 4.2
XLF Financials 4.0
XLV Health Care 3.9
XLI Industrials 3.7
XLB Materials 3.4
XLK Technology 4.7
XLU Utilities 3.8
Figure 2: The CIDR of three pairs of ETFs (1.XLF and XLK, 2.XLY and XLU, 3.XLY and XLE). For each pair, the curves representing the four most extreme days are displayed, with matching colors and line types indicating curves from the same day.
Figure 3: Estimates of the pairwise extremal correlation coefficients of CIDRs across the nine sectors.
We follow the step-by-step guide for estimating $\rho_{XY}$ presented in Section 5. First, for each sector, we center the curves around their sample mean functions, $\bar{R}_{N}(t)=\frac{1}{N}\sum_{i=1}^{N}R_{i}(t)$, and compute its norm $\|R_{i}(t)-\bar{R}_{N}(t)\|$ using (15) with $J=390$. We then examine whether the Hill plots of the norms for each sector exhibit stable regions. As shown in Fig 9 of Supplementary material, the norms appear regularly varying, so we compute the tail estimates $\hat{\alpha}$ for each sector using the Hill estimator, as shown in Table 3. Since the sectors are not tail equivalent, we apply the power transformation (16) to achieve tail equivalence with $\alpha=3$. This choice yields small biases in finite samples, as shown in Section 5. Next, we use the methods from Danielsson et al. (2016) to determine the optimal $k$ for estimating $\rho_{XY}$ for each pair across the sectors.
Figure 3 shows estimates of the pairwise extremal correlation coefficient across the nine ETF sectors. All pairs exhibit positive extremal correlations ($\hat{\rho}_{n,k}=0.43\sim 0.96$), and 44% of the pairs have strong extremal correlations above 0.7. We see that the CIDRs overall exhibit matching patterns of cumulative intraday returns on extreme market volatility days during the COVID-19 market turbulence, where most sectors either drop or increase together. However, our coefficient reveals more subtle information as well. For example, extreme return curves of XLF (Financials) are exceptionally strongly correlated with extreme curves for XLV, XLB and XLK (Health Care, Materials, Technology), but moderately correlated with XLU (Utilities). We do not aim at an analysis of the stock market or the economy, but we note that some findings are interesting. One might expect that the financial sector (mostly banks) to be strongly affected by the technology sector (mostly large IT companies like Google or Microsoft) because such mega corporations dominate the U.S. stock market. The similarity of extreme return curves for XLF and XLK is shown in the leftmost panels of Figure 2. One might expect bank stocks to be less affected by utility companies, whose revenues are largely fixed, but it’s less obvious why banks are strongly correlated with Health Care and Materials sectors. As another comparison, XLY (Consumer Discretionary) and XLU (Utilities) show a moderate extremal correlation of 0.43. Their extreme curves exhibit relatively dissimilar patterns, as seen in the middle panels in Figure 2.
Figure 4: Estimates of the pairwise coefficients of CIDRs, calculated from closing returns (left) and from all curves including non-extreme parts (right), are displayed.
We conclude by emphasizing that our tool offers a more precise quantification of intraday risk during extreme events. First, by analyzing curve shapes, it provides a better assessment of intraday risks. The left plot of Figure 4 displays the pairwise coefficients from single-valued closing returns. This plot reveals somewhat different information from Figure 3. For instance, between XLY and XLE, the closing returns show a $-0.3$ correlation, indicating negative correlation during extreme days. However, this value does not accurately reflect the positive relationship observed in the rightmost plots in Figure 2, where the paired extreme curves appear somewhat similar. Second, by focusing on the extreme parts of paired samples of curves, our tool effectively quantifies risk during these critical events. The right plot of Figure 4 shows correlation coefficients computed from all curves, including non-extreme ones. These coefficients tend to underestimate the extreme risk, highlighting the necessity for tools that specifically describe extreme conditions.
6.2 Extremal correlation between daily temperature curves
Figure 5: The three locations in the United States: Fort Collins, CO; Colorado Springs, CO; Austin, TX. The pairwise extremal correlation of daily temperature curves between the three locations is evaluated.
In this section, we evaluate the tendency of paired daily temperature curves to exhibit similar extreme patterns across three locations in the United States. The three locations are marked in Figure 5. We focus on the pairwise extremal dependence of those curves during the 2021 heat wave. Although this example focuses on temperature curves, our tool can be used for analyzing other curves during extreme weather events; for example, daily precipitation patterns or river flows during floods. A correlation of extreme data during past events may help with planning a resilient infrastructure that can better withstand the next extreme weather event.
We use hourly temperature measurements provided by the European Centre for Medium-Range Weather Forecasts (ECMWF). The data are part of their ERA5 (Fifth Generation of ECMWF atmospheric reanalyses) dataset, and represent the temperatures of air at 2 meters above the surface of land, sea or inland waters. We refer to Hersbach et al. (2020) for more details on the ERA5 data. We partition the hourly data into daily curves, with each day’s curve starting at midnight local time, to produce comparable daily temperature curves across locations in different time zones. We denote the temperature (in Celsius) on day $i$ at hour $t$ by $X_{i}(t)$, $i=1,\ldots,N$. Figure 6 depicts examples of daily temperature curves at the three locations. The data are taken from May 12, 2021 to Aug 31, 2021 ($N$ = 112).
Figure 6: Extreme daily temperature curves (in Celsius) during the 2021 heat wave (local time on the x-axis). Curves of matching color represent the same days when both Fort Collins and Colorado Springs experienced extreme patterns simultaneously.
We follow the step-by-step guide outlined in Section 5 to compute $\hat{\rho}_{n,k}$ for each pair of the three locations. First, for each location, daily curves are centered by the mean function, $\bar{X}_{N}(t)=\frac{1}{N}\sum_{i=1}^{N}X_{i}(t)$, and its norm $\|X_{i}(t)-\bar{X}_{N}(t)\|$ is computed using (15) with $J=24$. We then examine whether the Hill plots of the norms for each location show stable regions. As seen in Figure 10 of Supplementary material, the norms are regularly varying, so we compute the tail estimates $\hat{\alpha}$ using the Hill estimator, as shown in Table 3. Since the marginals across the three locations are not tail-equivalent, we apply the power transformation (16) to achieve tail equivalence with $\alpha=3$. We then apply the method from Danielsson et al. (2016) to determine the optimal $k$ for estimating $\rho_{XY}$ for each pair.
Table 4: Tail index estimates $\hat{\alpha}$ and pairwise extremal correlation coefficients $\hat{\rho}_{n,k}$ of daily temperature curves across Fort Collins, CO, Colorado Springs, CO, and Austin, TX.
$\hat{\rho}_{n,k}$
Location $\hat{\alpha}$ Fort Colorado Austin
Collins Springs
Fort Collins 4.4 1 0.98 0.83
Colorado Springs 3.8 0.98 1 0.86
Austin 3.4 0.83 0.86 1
Table 4 reports estimates of the pairwise extremal correlation coefficient across the three locations. There are positive and strong extremal correlations among all pairs ($\hat{\rho}_{n,k}=0.83\sim 0.98$), suggesting a high degree of association between the daily temperature extreme patterns across the three locations, even between different climatic regions like the Front Range foothills and the southern edge of the Great Plains. We see that the proximity in geographical locations corresponds to greater similarity in extreme patterns, showing that $\hat{\rho}_{n,k}$ is a meaningful and useful dependence measure.
7 Discussion
We introduced a coefficient designed to measure extremal dependence in paired samples of functions. This coefficient specifically focuses on extreme curves and computes their inner product to provide a quantitative assessment of the risk that paired extreme curves will occur simultaneously. The new coefficient is based on the theory of regular variation in $L^{2}$, providing a solid foundation for its derivation. Its estimator is shown to be consistent, a result supported by simulations. Deriving its asymptotic distribution might lead to a method for constructing confidence intervals. On the practical side, the estimator’s performance may depend on the method used to select an optimal $k$, as discussed in Section 5. While we considered two methods based on KS distances, that produce satisfactory results, future work could explore alternative methods and refine the selection of $k$.
Despite promising results, there are some limitation to our approach that suggest potential directions for future work. We treat functional observations as regularly varying square integrable random functions in $L^{2}$, which requires $\alpha>2$. While this condition is met for the financial and environmental data we work with, it might be desirable to find a correlation-like extremal dependence measure that requires merely $\alpha>0$. Possible extensions could involve the codifference or the covariation introduced in Kokoszka & Taqqu (1995). These measures of dependence are applicable to stable vectors with the index $\alpha<2$, and have been studied in econometrics and statistical physics contexts, see e.g., Kokoszka & Taqqu (1996) and Levy & Taqqu (2014) and Wolymańska et al. (2015). Exploring extreme value theory for functional data in this context might be useful, but one must note that those measures are not symmetric.
Acknowledgement
We Thank Professor Hong Miao of the Department of Finance and Real Estate at Colorado State University for preprocessing the financial data used in Section 6.1. We thank Professor Joshua French of the Department of Mathematical and Statistical Sciences at the University of Colorado Denver for preprocessing the temperature data used in Section 6.2. Kim and Kokoszka were partially supported by the United States National Science Foundation.
Supplementary material
The Supplementary material contains proofs of the theoretical results stated in the main paper and additional empirical results.
Supplementary material for “Extremal correlation coefficient for functional data”
[…]
Appendix A Preliminary results
In this section, we put together preliminary results needed to prove Theorem 4.4. Some of these results are known in the literature, and none of them are particularly profound or difficult to prove. However, these results allow us to streamline the exposition of proofs of the main result. Recall that, c.f., (5), $\mathcal{A}_{r}=\{(x,y)\in{\mathbb{B}}_{0}^{2}:\|(x,y)\|_{{\mathbb{B}}^{2}}\geq r\}$, $r>0$, where $\|(x,y)\|_{{\mathbb{B}}^{2}}=\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}$.
Lemma A.1.
Suppose $\mu$ is a measure in $M_{0}({\mathbb{B}}^{2})$ satisfying $\mu(t\cdot)=t^{-\alpha}\mu(\cdot)$, $t>0$. Then, $\mathcal{A}_{r}$ is a $\mu$–continuity set, i.e., $\mu(\partial\mathcal{A}_{r})=0$.
Proof A.2.
We assume $\mu(\partial{\mathcal{A}_{r}})>0$ and get a contradiction. Since ${\mathcal{A}_{r}}\supset\bigcup_{n\geq 1}\partial(n^{1/\alpha}{\mathcal{A}_{r}})$, it follows from the homogeneity property of $\mu$ that
$\mu({\mathcal{A}_{r}})\geq\sum_{n=1}^{\infty}\mu(\partial(n^{1/\alpha}{\mathcal{A}_{r}}))=\sum_{n=1}^{\infty}\mu(n^{1/\alpha}\partial{\mathcal{A}_{r}})=\sum_{n=1}^{\infty}n^{-1}\mu(\partial{\mathcal{A}_{r}})=\infty.$
It contradicts to the fact that $\mu$ is boundedly finite.
Recall that $R=\|(X,Y)\|$, $R_{i}=\|(X_{i},Y_{i})\|$, and $R_{(k)}$ is the $k$th largest order statistic with the convention $R_{(1)}=\max\{R_{1},\ldots,R_{n}\}$. Let $b(n)$ be the quantile function such that ${\rm pr}(R>b(n))=n^{-1}$. Then, the following lemma holds by Proposition 3.1 of Segers et al. (2017), and Theorem 4.1 and the proof of Theorem 4.2 of Resnick (2007).
Lemma A.3.
Let $M_{+}(0,\infty]$ be the space of Radon measures on $(0,\infty]$, and $\nu_{\alpha}(r,\infty]=r^{-\alpha}$. Also, let $\epsilon_{x}(A)=1$ if $x\in A$ and $\epsilon_{x}(A)=0$ if $x\notin A$. If $[X,Y]^{\top}$ is regularly varying in $L^{2}\times L^{2}$ according to Definition 3.1, then
(i) $R$ is a nonnegative random variable whose distribution has a regularly varying tail with index $-\alpha$,
(ii) $\frac{1}{k}\sum_{i=1}^{n}\epsilon_{R_{i}/b(n/k)}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\nu_{\alpha}$, in $M_{+}(0,\infty]$,
(iii) $R_{(k)}/b(n/k)\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}1$, in $[0,\infty)$,
(iv) $\frac{1}{k}\sum_{i=1}^{n}\epsilon_{R_{i}/R_{(k)}}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\nu_{\alpha}$ in $M_{+}(0,\infty]$.
The following lemma is used to prove Lemmas A.6 and A.8.
Lemma A.4.
Suppose $\gamma_{n}$ converges vaguely to $\nu_{\alpha}$ in $M_{+}(0,\infty]$. Then for any compact interval $K\subset(0,\infty]$,
$\int_{K}r^{2}\gamma_{n}(dr)\to\int_{K}r^{2}\nu_{\alpha}(dr).$
Proof A.5.
Since the function $r\mapsto r^{2}I_{K}$ is not continuous, we use an approximation argument. Set $K=[a,b]$, for $0<a<b\leq\infty$. Construct compact intervals $K_{j}\searrow K$ and nonnegative continuous functions $f_{j}$ such that $I_{K}\leq f_{j}\leq I_{K_{j}}$. By the triangle inequality,
$\left|\int_{K}r^{2}\gamma_{n}(dr)-\int_{K}r^{2}\nu_{\alpha}(dr)\right|$ $\leq\left|\int r^{2}I_{K}(r)\gamma_{n}(dr)-\int r^{2}f_{j}(r)\gamma_{n}(dr)\right|$
$\ \ +\left|\int r^{2}f_{j}(r)\gamma_{n}(dr)-\int r^{2}f_{j}(r)\nu_{\alpha}(dr)\right|$
$\ \ +\left|\int r^{2}f_{j}(r)\nu_{\alpha}(dr)-\int r^{2}I_{K}(r)\nu_{\alpha}(dr)\right|$
$=:A_{n,j}^{(1)}+A_{n,j}^{(2)}+A_{j}^{(3)}.$
Fix $\tau>0$. There is $j^{\star}$ such that for $j\geq j^{\star}$,
$A_{j}^{(3)}\leq c\int\left[f_{j}(r)-I_{K}(r)\right]\nu_{\alpha}(dr)\leq c\nu_{\alpha}(K_{j}\setminus K^{\circ})<\tau/2,$
where $c=b^{2}I_{b\neq\infty}+a^{2}I_{b=\infty}$. Similarly $A_{n,j}^{(1)}\leq c\gamma_{n}(K_{j}\setminus K^{\circ})$, so for every fixed $j$,
$\limsup_{n\to\infty}A_{n,j}^{(1)}\leq M^{2}\limsup_{n\to\infty}\gamma_{n}(K_{j}\setminus K^{\circ})\leq M^{2}\nu_{\alpha}(K_{j}\setminus K^{\circ})$
because $K_{j}\setminus K^{\circ}$ is compact, cf. Proposition 3.12 in Resnick (1987). Thus,
$\limsup_{n\to\infty}\left|\int_{K}r^{2}\gamma_{n}(dr)-\int_{K}r^{2}\nu_{\alpha}(dr)\right|\leq\tau+\limsup_{n\to\infty}A_{n,j^{\star}}^{(2)}=\tau.$
Since $\tau$ is arbitrary, we get the claim.
The following two lemmas are used to prove Lemma A.10 and Proposition B.1.
Lemma A.6.
Under Assumption 4, for any $M>0$,
$\frac{n}{k}E\left[\left(\frac{R}{b(n/k)}\right)^{2}I_{R\geq Mb(n/k)}\right]\to\frac{\alpha}{\alpha-2}M^{2-\alpha}.$
Proof A.7.
Observe that
$\frac{n}{k}E\left[\left(\frac{R}{b(n/k)}\right)^{2}I_{R\geq Mb(n/k)}\right]=\int_{M}^{\infty}r^{2}\frac{n}{k}{\rm pr}\left(\frac{R}{b(n/k)}\in dr\right),$
and
$\frac{\alpha}{\alpha-2}M^{2-\alpha}=\int_{M}^{\infty}r^{2}\nu_{\alpha}(dr).$
By Lemma A.3 (i), we have that in $M_{+}(0,\infty]$
$\frac{n}{k}{\rm pr}\left(\frac{R}{b(n/k)}\in\cdot\right)\stackrel{{\scriptstyle\mbox{$\scriptstyle v$}}}{{\rightarrow}}\nu_{\alpha}.$
Therefore, we get the claim by Lemma A.4 with $K=[M,\infty]$.
Lemma A.8.
The function $h$ on $M_{+}(0,\infty]$ defined by $h(\gamma)=\int_{1}^{M}r^{2}\gamma(dr)$ is continuous at $\nu_{\alpha}$.
Proof A.9.
Suppose $\gamma_{n}$ converges vaguely to $\nu_{\alpha}$. Then, by Lemma A.4 with $K=[1,M]$, it can be shown that
$\lim_{n\to\infty}\int_{1}^{M}r^{2}\gamma_{n}(dr)=\int_{1}^{M}r^{2}\nu_{\alpha}(dr).$
The following lemma is the key argument to prove Proposition B.3.
Lemma A.10.
Under Assumption 4, the following statements hold:
[…]
Proof A.11.
The proofs for (18) and (19) are almost the same, so we only prove (18) to save space. Let $\hat{\gamma}_{n,k}=\frac{1}{k}\sum_{i=1}^{n}\epsilon_{R_{i}/R_{(k)}}$, and recall that $\hat{\gamma}_{n,k}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\nu_{\alpha}$ (see Lemma A.3 (iv)). Since
$\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{R_{(k)}}\right)^{2}I_{R_{i}\geq R_{(k)}}=\int_{1}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr),$
we need to show that
$\int_{1}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr)\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\int_{1}^{\infty}r^{2}\nu_{\alpha}(dr)=\frac{\alpha}{\alpha-2}.$
To prove this convergence, we use the second converging together theorem, Theorem 3.5 in Resnick (2007), (also stated as Theorem 3.2 of Billingsley (1999)).
Let
$V_{n,k}=\int_{1}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr),$ $V=\int_{1}^{\infty}r^{2}\nu_{\alpha}(dr);$
$V_{n,k}^{(M)}=\int_{1}^{M}r^{2}\hat{\gamma}_{n,k}(dr),$ $V^{(M)}=\int_{1}^{M}r^{2}\nu_{\alpha}(dr).$
To show the desired convergence $V_{n,k}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}V$ (equivalently, $V_{n,k}\stackrel{{\scriptstyle\mbox{$\scriptstyle d$}}}{{\rightarrow}}V$), we must verify that
$\forall\ M>1,\ \ \ V_{n,k}^{(M)}\stackrel{{\scriptstyle\mbox{$\scriptstyle d$}}}{{\rightarrow}}V^{(M)},\ \ \ \ {\rm as}\ n\to\infty;$ (20)
$V^{(M)}\stackrel{{\scriptstyle\mbox{$\scriptstyle d$}}}{{\rightarrow}}V,\ \ \ {\rm as}\ M\to\infty;$ (21)
$\forall\ \varepsilon>0,\ \ \ \lim_{M\to\infty}\limsup_{n\to\infty}{\rm pr}\left(|V_{n,k}^{(M)}-V_{n,k}|>\varepsilon\right)=0.$ (22)
Convergence (20) follows from Lemma A.3 (iv) and Lemma A.8. Convergence (21) holds since for $\alpha>2$
[…]
It remains to show that $\forall\varepsilon>0$,
$\lim_{M\to\infty}\limsup_{n\to\infty}{\rm pr}\left(|V_{n,k}^{(M)}-V_{n,k}|>\varepsilon\right)=\lim_{M\to\infty}\limsup_{n\to\infty}{\rm pr}\left(\int_{M}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr)>\varepsilon\right)=0.$
Fix $\varepsilon>0$ and $\eta>0$. Observe that
${\rm pr}\left(\int_{M}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr)>\varepsilon\right)\leq Q_{1}(n)+Q_{2}(n),$
where
$Q_{1}(n)={\rm pr}\left(\int_{M}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr)>\varepsilon,\ \left|\frac{R_{(k)}}{b(n/k)}-1\right|<\eta\right),\ \ Q_{2}(n)={\rm pr}\left(\left|\frac{R_{(k)}}{b(n/k)}-1\right|\geq\eta\right).$
By Lemma A.3 (iii), $\limsup_{n\to\infty}Q_{2}(n)=0$. For $Q_{1}(n)$, we start with the bound
$Q_{1}(n)$ $\leq{\rm pr}\left(\int_{M}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr)>\varepsilon,\ \frac{R_{(k)}}{b(n/k)}>1-\eta\right)$
$={\rm pr}\left(\int_{M}^{\infty}r^{2}\frac{1}{k}\sum_{i=1}^{n}\epsilon_{R_{i}/R_{(k)}}(dr)>\varepsilon,\ \frac{R_{(k)}}{b(n/k)}>1-\eta\right).$
Conditions $R_{i}/R_{(k)}>M$ and ${R_{(k)}}/{b(n/k)}>1-\eta$ imply ${R_{i}}/{b(n/k)}>M(1-\eta)$, so
$Q_{1}(n)$ $\leq{\rm pr}\left(\int_{M(1-\eta)}^{\infty}r^{2}\frac{1}{k}\sum_{i=1}^{n}\epsilon_{R_{i}/b(n/k)}(dr)>\varepsilon\right)$
$={\rm pr}\left(\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq M(1-\eta)b(n/k)}>\varepsilon\right).$
Then, it follows from Markov’s inequality and Lemma A.6 that
$Q_{1}(n)\leq\frac{1}{\varepsilon}\frac{n}{k}E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq M(1-\eta)b(n/k)}\right]\to\frac{1}{\varepsilon}\frac{\alpha}{\alpha-2}\{M(1-\eta)\}^{2-\alpha},\ \ \ {\rm as}\ n\to\infty.$
This bound goes to 0 as $M\to\infty$ since $\alpha>2$.
The following lemma follows from Theorem 3.8 of McDiarmid (1998). It states a Bernstein type inequality, which is the key technique to prove Proposition B.1.
Lemma A.12.
Let ${\bf Z}_{n}=(Z_{1},\ldots,Z_{n})$ with the $Z_{i}$ taking values in a Lebesgue measurable subset ${\mathcal{Z}}$ of an Euclidean space. Let $f$ be a real-valued function defined on ${\mathcal{Z}}^{n}$. For $(z_{1},\ldots,z_{i})\in{\mathcal{Z}}^{i}$, $1\leq i\leq n$, put
$g_{i}(z_{1},\ldots,z_{i}):=E\left[f({\bf Z}_{n})|Z_{j}=z_{j},1\leq j\leq i\right]-E\left[f({\bf Z}_{n})|Z_{j}=z_{j},1\leq j\leq i-1\right].$ (23)
Define the maximum deviation by
$b:=\max_{1\leq i\leq n}\sup_{(z_{1},\ldots,z_{i})\in{\mathcal{Z}}^{i}}g_{i}(z_{1},\ldots,z_{i}),$ (24)
and define the supremum sum of variances by
$\hat{v}:=\sup_{(z_{1},\ldots,z_{n})\in{\mathcal{Z}}^{n}}\sum_{i=1}^{n}\hbox{\rm var}\left[g_{i}(z_{1},\ldots,z_{i-1},Z_{i}^{\prime})\right],$ (25)
where $Z_{i}^{\prime}$ is an independent copy of $Z_{i}$ conditional on $Z_{j}=z_{j}$, $11\leq j\leq i-$ 1. If $b$ and $\hat{v}$ are finite, then for any $\varepsilon\geq 0$,
${\rm pr}\left(f({\bf Z}_{n})-E[f({\bf Z}_{n})]\geq t\right)\leq\exp\left(\frac{-\varepsilon^{2}}{2(\hat{v}+b\varepsilon/3)}\right).$
[…]
Appendix B Proof of Theorem 4.4 in Section 4
Recall (13), i.e., the definition:
$\hat{\sigma}_{n,k}=\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{R_{(k)}},\frac{Y_{i}}{R_{(k)}}\right\rangle I_{R_{i}\geq R_{(k)}}.$
To prove the consistency of $\hat{\sigma}_{n,k}$ for the extremal covariance $\sigma_{XY}$, we consider the following sequence of random variables
$\sigma_{n,k}:=\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{b(n/k)},\frac{Y_{i}}{b(n/k)}\right\rangle I_{R_{i}\geq b(n/k)}.$ (26)
Note that $\sigma_{n,k}$ is not observable since $b(\cdot)$ is unknown. However, $b(n/k)$ can be estimated by its consistent estimator $R_{(k)}$, and it can be shown that replacing $b(n/k)$ by $R_{(k)}$ ensures that the difference between $\sigma_{n,k}$ and $\hat{\sigma}_{n,k}$ is asymptotically negligible, which will be shown in Proposition B.3. Thus, the key argument for establishing the consistency is to show that $\sigma_{n,k}$ converges in probability to $\sigma_{XY}$, which is proven in the following proposition.
Proposition B.1.
Under Assumption 4,
$\sigma_{n,k}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\sigma_{XY}.$
Proof B.2.
Set
$\bar{\sigma}_{n,k}:=E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle\Bigg{|}\|X_{1}\|\vee\|Y_{1}\|>b(n/k)\right].$ (27)
Then, by Proposition 4.2, $\bar{\sigma}_{n,k}\to\sigma_{XY}$, so it remains to show that $|\sigma_{n,k}-\bar{\sigma}_{n,k}|\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0$.
Let ${\bf Z}_{n}=(Z_{1},\ldots,Z_{n})$, where $Z_{i}=(X_{i},Y_{i})$, and ${\bf z}_{n}=(z_{1},\ldots,z_{n})$, where $z_{i}=(x_{i},y_{i})$, for $1\leq i\leq n$. Consider a map $f:(L^{2}\times L^{2})^{n}\to{\mathbb{R}}$ defined by
$f({\bf z}_{n}):=\left|\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{x_{i}}{b(n/k)},\frac{y_{i}}{b(n/k)}\right\rangle I_{r_{i}\geq b(n/k)}-\frac{n}{k}E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}>b(n/k)}\right]\right|.$
Then, we have that
$|\sigma_{n,k}-\bar{\sigma}_{n,k}|=f({\bf Z}_{n})-E[f({\bf Z}_{n})]+E[f({\bf Z}_{n})].$
We aim to show that $f({\bf Z}_{n})-E[f({\bf Z}_{n})]\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0$ and $E[f({\bf Z}_{n})]\to 0$.
To establish the convergence, $f({\bf Z}_{n})-E[f({\bf Z}_{n})]\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0$, we use the Bernstein type concentration inequality in Lemma A.12. Since the $(X_{i},Y_{i})$ are independent, the deviation function in (23) has the following form
$g_{i}(z_{1},\ldots,z_{i})=E\left[f(z_{1},\ldots,z_{i-1},z_{i},Z_{i+1},\ldots,Z_{n})-f(z_{1},\ldots,z_{i-1},Z_{i},Z_{i+1},\ldots,Z_{n})\right].$
Then, using the fact that $\left||x|-|y|\right|\leq|x-y|$, we have that
$g_{i}(z_{1},\ldots,z_{i})$ $\leq\frac{1}{k}E\left[\left|\left\langle\frac{x_{i}}{b(n/k)},\frac{y_{i}}{b(n/k)}\right\rangle I_{r_{i}\geq b(n/k)}-\left\langle\frac{X_{i}}{b(n/k)},\frac{Y_{i}}{b(n/k)}\right\rangle I_{R_{i}\geq b(n/k)}\right|\right]$
$\leq\frac{1}{k}\left\{\frac{|\left\langle x_{i},y_{i}\right\rangle|}{b(n/k)^{2}}+\frac{k}{n}\frac{n}{k}E\left[\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\right]\right\}$
$\leq\frac{1}{k}\left\{\frac{\|x_{i}\|\|y_{i}\|}{b(n/k)^{2}}+\frac{n}{k}E\left[\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\right]\right\}.$
Since $(x_{i},y_{i})\in L^{2}\times L^{2}$ and $\frac{n}{k}E\left[\left({R_{i}}/{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\right]\to\alpha/(\alpha-2)$ by Lemma A.6, we have that $g_{i}(z_{1},\ldots,z_{i})\leq{c_{1}}/{k}$, for some constant $c_{1}>0$. Therefore, the maximum deviation $b$ in (24) is bounded by $c_{1}/k$.
Next we investigate the upper bound for the sum of variances $\hat{v}$ in (25). Since $E[g_{i}(z_{1},\ldots,z_{i-1},Z_{i}^{\prime})]=0$ by the law of total probability, we have that
$\hbox{\rm var}\left[g_{i}(z_{1},\ldots,z_{i-1},Z_{i}^{\prime})\right]$
$=E[g_{i}^{2}(z_{1},\ldots,z_{i-1},Z_{i}^{\prime})]$
$=E\left[\left\{f(z_{1},\ldots,z_{i-1},Z_{i}^{\prime},Z_{i+1},\ldots,Z_{n})-f(z_{1},\ldots,z_{i-1},Z_{i},Z_{i+1},\ldots,Z_{n})\right\}^{2}\right]$
$\leq\frac{1}{k^{2}}E\left[\left\{\left\langle\frac{X_{i}^{\prime}}{b(n/k)},\frac{Y_{i}^{\prime}}{b(n/k)}\right\rangle I_{R_{i}^{\prime}\geq b(n/k)}-\left\langle\frac{X_{i}}{b(n/k)},\frac{Y_{i}}{b(n/k)}\right\rangle I_{R_{i}\geq b(n/k)}\right\}^{2}\right]$
$\leq\frac{2}{k^{2}}E\left[\left\langle\frac{X_{i}}{b(n/k)},\frac{Y_{i}}{b(n/k)}\right\rangle^{2}I_{R_{i}\geq b(n/k)}\right]$
$\leq\frac{2}{k^{2}}\left\{\frac{k}{n}\frac{n}{k}E\left[\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\right]\right\}.$
It then again follows from Lemma A.6 that $\hbox{\rm var}\left[g_{i}(z_{1},\ldots,z_{i-1},Z_{i}^{\prime})\right]\leq c_{2}/(nk)$ for some $c_{2}>0$. Then the supremum sum of variances $\hat{v}$ is bounded above by $c_{2}/k$. Therefore by Lemma A.12, for any $\varepsilon>0$
${\rm pr}\left(f({\bf Z}_{n})-E[f({\bf Z}_{n})]\geq\varepsilon\right)\leq\exp\left(\frac{-k\varepsilon^{2}}{c_{1}+c_{2}\varepsilon/3}\right).$
If we apply this inequality to $-f({\bf Z}_{n})$, then we obtain the following ‘two-sided’ inequality
${\rm pr}\left(|f({\bf Z}_{n})-E[f({\bf Z}_{n})]|\geq\varepsilon\right)\leq 2\exp\left(\frac{-k\varepsilon^{2}}{c_{1}+c_{2}\varepsilon/3}\right).$
From this, we obtain that $f({\bf Z}_{n})-E[f({\bf Z}_{n})]\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0$.
Next, to show $E[f({\bf Z}_{n})]\to 0$, we set, for $1\leq i\leq n$
$\Delta_{i}=\left\langle\frac{X_{i}}{b(n/k)},\frac{Y_{i}}{b(n/k)}\right\rangle I_{R_{i}\geq b(n/k)}-E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}>b(n/k)}\right].$
Then, we have that
$E\left[f({\bf Z}_{n})\right]=\frac{n}{k}E\left[\left|\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}\right|\right]$ $\leq\frac{n}{k}\left\{E\left[\left(\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}\right)^{2}\right]\right\}^{1/2}$
$=\frac{n}{k}\left\{E\left[\frac{1}{n^{2}}\sum_{i=1}^{n}\Delta_{i}^{2}+\frac{1}{n^{2}}\sum_{i\neq j}\Delta_{i}\Delta_{j}\right]\right\}^{1/2}.$
Since the $\Delta_{i}$ are independent, $E[\Delta_{i}\Delta_{j}]=0$, for $i\neq j$. Therefore,
$E\left[f({\bf Z}_{n})\right]$
$\leq\frac{\sqrt{n}}{k}\left\{E\left[\Delta_{1}^{2}\right]\right\}^{1/2}$
$=\frac{\sqrt{n}}{k}\left\{E\left[\left(\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}\geq b(n/k)}-E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}>b(n/k)}\right]\right)^{2}\right]\right\}^{1/2}$
$=\frac{\sqrt{n}}{k}\left\{\hbox{\rm var}\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}$
$\leq\frac{\sqrt{n}}{k}\left\{E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}$
$\leq\frac{\sqrt{n}}{k}\left\{E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}.$
Therefore, by Lemma A.6 we have that
$E\left[f({\bf Z}_{n})\right]\leq\frac{\sqrt{n}}{k}\left\{\frac{k}{n}\frac{n}{k}E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}\leq\frac{c_{3}}{\sqrt{k}},$
for some $c_{3}>0$, which completes the proof.
Proposition B.3.
Under Assumption 4,
$|\hat{\sigma}_{n,k}-\sigma_{n,k}|\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0.$
Proof B.4.
Consider the following decomposition
$|\hat{\sigma}_{n,k}-\sigma_{n,k}|\leq P_{1}(n)+P_{2}(n),$
where
$P_{1}(n):=\left|\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{R_{(k)}},\frac{Y_{i}}{R_{(k)}}\right\rangle\left\{I_{R_{i}\geq R_{(k)}}-I_{R_{i}\geq b(n/k)}\right\}\right|,$
$P_{2}(n):=\left|\frac{1}{k}\sum_{i=1}^{n}\left\{\left\langle\frac{X_{i}}{R_{(k)}},\frac{Y_{i}}{R_{(k)}}\right\rangle-\left\langle\frac{X_{i}}{b(n/k)},\frac{Y_{i}}{b(n/k)}\right\rangle\right\}I_{R_{i}\geq b(n/k)}\right|.$
We will show that each of the two parts goes to 0. We first focus on $P_{1}(n)$. Observe that
$P_{1}(n)$ $\leq\left(\frac{b(n/k)}{R_{(k)}}\right)^{2}\frac{1}{k}\sum_{i=1}^{n}\left|\left\langle\frac{X_{i}}{R_{i}},\frac{Y_{i}}{R_{i}}\right\rangle\right|\left(\frac{R_{i}}{b(n/k)}\right)^{2}\left|I_{R_{i}\geq R_{(k)}}-I_{R_{i}\geq b(n/k)}\right|$
$\leq\left(\frac{b(n/k)}{R_{(k)}}\right)^{2}\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}\left|I_{R_{i}\geq R_{(k)}}-I_{R_{i}\geq b(n/k)}\right|$
$=\left(\frac{b(n/k)}{R_{(k)}}\right)^{2}\left|\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq R_{(k)}}-\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\right|$
$=\left|\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{R_{(k)}}\right)^{2}I_{R_{i}\geq R_{(k)}}-\left(\frac{b(n/k)}{R_{(k)}}\right)^{2}\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\right|$
Then, by Lemma A.3 (iii), we have that $\left({b(n/k)}/{R_{(k)}}\right)^{2}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}1$. By Lemma A.10 that $\frac{1}{k}\sum_{i=1}^{n}\left({R_{i}}/{R_{(k)}}\right)^{2}I_{R_{i}\geq R_{(k)}}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\alpha/(\alpha-2)$ and $\frac{1}{k}\sum_{i=1}^{n}\left({R_{i}}/{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\alpha/(\alpha-2)$. Therefore, we have that $P_{1}(n)\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0$.
Now we work on $P_{2}(n)$. Observe that
$P_{2}(n)$ $=\left|\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{R_{i}},\frac{Y_{i}}{R_{i}}\right\rangle R_{i}^{2}\left(\frac{1}{R_{(k)}^{2}}-\frac{1}{b(n/k)^{2}}\right)I_{R_{i}\geq b(n/k)}\right|$
$\leq\left|\frac{b(n/k)^{2}}{R_{(k)}^{2}}-1\right|\frac{1}{k}\sum_{i=1}^{n}\left|\left\langle\frac{X_{i}}{R_{i}},\frac{Y_{i}}{R_{i}}\right\rangle\right|\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}$
$\leq\left|\frac{b(n/k)^{2}}{R_{(k)}^{2}}-1\right|\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}.$
By Lemma A.6, we have that $\frac{1}{k}\sum_{i=1}^{n}\left({R_{i}}/{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}=O_{P}(1)$, and by Lemma A.3 (iii), we have that $b(n/k)/R_{(k)}\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}1$. Thus, $P_{2}(n)\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}0$.
Proof of Theorem 4.4. It follows from Propositions B.1 and B.3.
[…]
Appendix C Proof of Lemma 5.1 in Section 5 and its extension to randomly sampled weights
Proof of Lemma 5.1:  We begin by noting that since $Z_{1}$ and $Z_{2}$ are independent, there exists $\nu$ in $M_{+}({\mathbb{R}}_{+}^{2})$ such that
$n{\rm pr}\left(\frac{(|Z_{1}|,|Z_{2}|)}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle\mbox{$\scriptstyle v$}}}{{\rightarrow}}\nu,$ (28)
and for ${\bf x}=[x_{1},x_{2}]^{\top}$
$\nu([0,{\bf x}]^{c})=c\{(x_{1})^{-\alpha}+(x_{2})^{-\alpha}\}.$
With the choice of $b(n)$ defined by
$n^{-1}$ $={\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))$ (29)
$={\rm pr}(|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)),$
we set $c=1/(1+(1-\rho^{2})^{\alpha/2})$ to ensure that $\nu$ is a probability measure on $\{(z_{1},z_{2}):|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1\}$.
We claim that
$\sigma_{XY}=\rho\frac{c\alpha}{\alpha-2};$ (30)
$\sigma_{X}^{2}=\frac{c\alpha}{\alpha-2};$ (31)
$\sigma_{Y}^{2}=\left\{\rho^{2}+(1-\rho^{2})^{\alpha/2}\right\}\frac{c\alpha}{\alpha-2}.$ (32)
We first work on (30). Since the terms with the $N_{j}$ do not affect the extremal behavior of $X$ and $Y$, we have that by Proposition 4.2
$\sigma_{XY}$
$=\lim_{n\to\infty}E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}}{b(n)}\right\rangle\Bigg{|}\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}\times$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}}{b(n)}\right\rangle I_{\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)}\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}E\left[\rho\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
It then follows from (28) and (29) that
$\sigma_{XY}$ $=\lim_{n\to\infty}nE\left[\rho\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
$=\lim_{n\to\infty}\int_{{\mathbb{R}}^{2}_{+}}\rho z_{1}^{2}I_{|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1}n{\rm pr}\left(\frac{|Z_{1}|}{b(n)}\in dz_{1},\frac{|Z_{2}|}{b(n)}\in dz_{2}\right)$
$=\int_{{\mathbb{R}}_{+}^{2}}\rho z_{1}^{2}I_{|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1}\ \nu(dz_{1},dz_{2})$
$=\int_{{\mathbb{R}}_{+}}\rho z_{1}^{2}I_{\{(z_{1},0):z_{1}>1\}}\ c\nu_{\alpha}(dz_{1})+\int_{{\mathbb{R}}_{+}}\rho z_{1}^{2}I_{\{(0,z_{2}):z_{2}>1/(1-\rho^{2})^{1/2}\}}\ c\nu_{\alpha}(dz_{2})$
$=\int_{1}^{\infty}\rho z_{1}^{2}\ c\nu_{\alpha}(dz_{1})+0=\rho\frac{c\alpha}{\alpha-2}.$
Analogously, for (31) we can show that
$\sigma_{X}^{2}$
$=\lim_{n\to\infty}E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{Z_{1}\phi_{1}}{b(n)}\right\rangle\Bigg{|}\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}E\left[\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
$=\lim_{n\to\infty}nE\left[\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
$=\frac{c\alpha}{\alpha-2}.$
Next, we work on (32). Observe that
$\sigma_{Y}^{2}$
$=\lim_{n\to\infty}E\left[\frac{\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|^{2}}{b(n)^{2}}\Bigg{|}\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}\times$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ E\left[\frac{\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right].$
Then, again it follows from (28) and (29) that
$\sigma_{Y}^{2}$ $=\lim_{n\to\infty}nE\left[\frac{\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
$=\lim_{n\to\infty}\int_{{\mathbb{R}}^{2}_{+}}\left\{\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2}\right\}I_{|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1}n{\rm pr}\left(\frac{|Z_{1}|}{b(n)}\in dz_{1},\frac{|Z_{2}|}{b(n)}\in dz_{2}\right)$
$=\int_{{\mathbb{R}}_{+}^{2}}\left\{\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2}\right\}I_{|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1}\ \nu(dz_{1},dz_{2})$
$=\int_{{\mathbb{R}}_{+}}\rho^{2}z_{1}^{2}I_{\{(z_{1},0):z_{1}>1\}}\ c\nu_{\alpha}(dz_{1})+\int_{{\mathbb{R}}_{+}}(1-\rho^{2})z_{2}^{2}I_{\{(0,z_{2}):z_{2}>1/(1-\rho^{2})^{1/2}\}}\ c\nu_{\alpha}(dz_{2})$
$=\int_{1}^{\infty}\rho^{2}z_{1}^{2}\ c\nu_{\alpha}(dz_{1})+\int_{1/(1-\rho^{2})^{1/2}}^{\infty}(1-\rho^{2})z_{2}^{2}\ c\nu_{\alpha}(dz_{1})$
=$=\left\{\rho^{2}+(1-\rho^{2})^{\alpha/2}\right\}\frac{c\alpha}{\alpha-2}.$
This completes the proof of Lemma 5.1.
We now extend Lemma 5.1 by considering randomly occurring weights, so each direction in the function space can contribute either a heavy-tailed or a light-weight component.
Lemma C.1.
Let $\{A_{i}\}$ and $\{B_{i}\}$ be independent sequences of iid Bernoulli random variables with $P(A_{i}=1)=p_{A}$ and $P(B_{i}=1)=p_{B}$, independent of the $Z_{i}$. Put
$X(t)=\sum_{i=1}^{2}\phi_{i}(t)\{Z_{i}A_{i}+N_{i}(1-A_{i})\},\quad Y(t)=\sum_{i=1}^{2}\phi_{i}(t)\{Z_{i}B_{i}+N_{i}(1-B_{i})\}.$
Then, under assumptions in Lemma 5.1, we have that
$\rho_{XY}=(p_{A})^{1/2}(p_{B})^{1/2}.$
Proof C.2.
We will show that
$\sigma_{X}^{2}=\frac{c\alpha}{\alpha-2}p_{A};\quad\sigma_{Y}^{2}=\frac{c\alpha}{\alpha-2}p_{B};\quad\sigma_{XY}=\frac{c\alpha}{\alpha-2}p_{A}p_{B}.$
We first work on $\sigma_{X}^{2}$. With the choice of $b(n)$ defined by $n^{-1}={\rm pr}((A_{1}^{2}Z_{1}^{2}+A_{2}^{2}Z_{2}^{2})^{1/2}\vee(B_{1}^{2}Z_{1}^{2}+B_{2}^{2}Z_{2}^{2})^{1/2}>b(n))$, it follows from the law of total expectation that
$\sigma_{X}^{2}$ $=\lim_{n\to\infty}nE\left[\frac{A_{1}^{2}Z_{1}^{2}+A_{2}^{2}Z_{2}^{2}}{b(n)^{2}}I_{(A_{1}^{2}Z_{1}^{2}+A_{2}^{2}Z_{2}^{2})^{1/2}\vee(B_{1}^{2}Z_{1}^{2}+B_{2}^{2}Z_{2}^{2})^{1/2}>b(n)}\right]$
$=\lim_{n\to\infty}nE\left[\frac{Z_{1}^{2}+Z_{2}^{2}}{b(n)^{2}}I_{(Z_{1}^{2}+Z_{2}^{2})^{1/2}\vee(B_{1}^{2}Z_{1}^{2}+B_{2}^{2}Z_{2}^{2})^{1/2}>b(n)}\right]\times p_{A}^{2}$
$+\lim_{n\to\infty}nE\left[\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(B_{1}^{2}Z_{1}^{2}+B_{2}^{2}Z_{2}^{2})^{1/2}>b(n)}\right]\times p_{A}(1-p_{A})$
$+\lim_{n\to\infty}nE\left[\frac{Z_{2}^{2}}{b(n)^{2}}I_{|Z_{2}|\vee(B_{1}^{2}Z_{1}^{2}+B_{2}^{2}Z_{2}^{2})^{1/2}>b(n)}\right]\times(1-p_{A})p_{A}.$
Then, using (28) and vague convergence, it simplifies to
$\sigma_{X}^{2}$ $=\left[\int_{{\mathbb{R}}_{+}}z_{1}^{2}I_{\{(z_{1},0):z_{1}>1\}}\ c\nu_{\alpha}(dz_{1})+\int_{{\mathbb{R}}_{+}}z_{2}^{2}I_{\{(0,z_{2}):z_{2}>1\}}\ c\nu_{\alpha}(dz_{2})\right]p_{A}^{2}$
$+\left[\int_{{\mathbb{R}}_{+}}z_{1}^{2}I_{\{(z_{1},0):z_{1}>1\}}\ c\nu_{\alpha}(dz_{1})\right]p_{A}(1-p_{A})+\left[\int_{{\mathbb{R}}_{+}}z_{2}^{2}I_{\{(z_{1},0):z_{2}>1\}}\ c\nu_{\alpha}(dz_{1})\right](1-p_{A})p_{A}$
$=\frac{2c\alpha}{\alpha-2}p_{A}^{2}+\frac{c\alpha}{\alpha-2}p_{A}(1-p_{A})+\frac{c\alpha}{\alpha-2}(1-p_{A})p_{A}=\frac{2c\alpha}{\alpha-2}p_{A}.$
Similarly, we can get $\sigma_{Y}^{2}=\frac{2c\alpha}{\alpha-2}p_{B}$.
Turning to $\sigma_{XY}$, we have that
$\sigma_{XY}$
$=\lim_{n\to\infty}nE\left[\frac{A_{1}B_{1}Z_{1}^{2}+A_{2}B_{2}Z_{2}^{2}}{b(n)^{2}}I_{(A_{1}^{2}Z_{1}^{2}+A_{2}^{2}Z_{2}^{2})^{1/2}\vee(B_{1}^{2}Z_{1}^{2}+B_{2}^{2}Z_{2}^{2})^{1/2}>b(n)}\right]$
$=\lim_{n\to\infty}nE\left[\frac{Z_{1}^{2}+Z_{2}^{2}}{b(n)^{2}}I_{(Z_{1}^{2}+Z_{2}^{2})^{1/2}>b(n)}\right]\times p_{A}^{2}p_{B}^{2}+\lim_{n\to\infty}nE\left[\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|>b(n)}\right]\times p_{A}p_{B}(1-p_{A}p_{B})$
$+\lim_{n\to\infty}nE\left[\frac{Z_{2}^{2}}{b(n)^{2}}I_{|Z_{2}|>b(n)}\right]\times(1-p_{A}p_{B})p_{A}p_{B}$
$=\left[\int_{{\mathbb{R}}_{+}}z_{1}^{2}I_{\{(z_{1},0):z_{1}>1\}}\ c\nu_{\alpha}(dz_{1})+\int_{{\mathbb{R}}_{+}}z_{2}^{2}I_{\{(0,z_{2}):z_{2}>1\}}\ c\nu_{\alpha}(dz_{2})\right]p_{A}^{2}p_{B}^{2}$
$+\left[\int_{{\mathbb{R}}_{+}}z_{1}^{2}I_{\{(z_{1},0):z_{1}>1\}}\ c\nu_{\alpha}(dz_{1})\right]p_{A}p_{B}(1-p_{A}p_{B})+\left[\int_{{\mathbb{R}}_{+}}z_{2}^{2}I_{\{(0,z_{2}):z_{2}>1\}}\ c\nu_{\alpha}(dz_{2})\right](1-p_{A}p_{B})p_{A}p_{B}$
$=\frac{2c\alpha}{\alpha-2}p_{A}^{2}p_{B}^{2}+\frac{c\alpha}{\alpha-2}p_{A}p_{B}(1-p_{A}p_{B})+\frac{c\alpha}{\alpha-2}(1-p_{A}p_{B})p_{A}p_{B}$
$=\frac{2c\alpha}{\alpha-2}p_{A}p_{B}.$
[…]
Appendix D Supplementary simulation results
D.1 Simulation results on the consistency of $\hat{\rho}_{n,k}$
This section reports the magnitude of empirical biases (the absolute difference between the average and the theoretical value), along with standard errors computed as the sample standard deviations. Using the optimal $k$s selected by the method from Danielsson et al. (2016), the results are shown in Tables 5 and 6 for $\alpha=\{4,5\}$. The results with the optimal $k$s selected by the method from Clauset et al. (2009) are provided in Tables 78, and 9, for $\alpha=\{3,4,5\}$.
In general, the estimators obtained with the method of Danielsson et al. (2016) exhibit substantially lower bias, but larger standard errors compared to those obtained with the method of Clauset et al. (2009). The lower bias is likely due to the fact that tail quantiles are particularly sensitive to small changes in probabilities. By minimizing the KS distance between the empirical and theoretical tail quantiles, as done by Danielsson et al. (2016), the method appears to achieve lower bias in finite samples. The larger standard errors result from this method selecting a much smaller value of $k$ compared to Clauset et al. (2009). In terms of MSE, no substantial difference appears to exist between the two methods.
Table 5: The magnitude of empirical biases (standard errors) of $\hat{\rho}_{n,k}$ when $\alpha=4$. Optimal $k$s are selected using the method from Danielsson et al. (2016), with averages of $k=9$ ($N=100$), $k=29$ ($N=500$), and $k=74$ ($N=2000$).
$\rho_{XY}$ $N=100$ $N=500$ $N=2000$
-1.0 0.08 (0.04) 0.06 (0.03) 0.04 (0.03)
-0.9 0.13 (0.11) 0.10 (0.09) 0.08 (0.08)
-0.8 0.12 (0.14) 0.10 (0.11) 0.07 (0.10)
-0.7 0.10 (0.15) 0.07 (0.12) 0.05 (0.11)
-0.6 0.07 (0.15) 0.06 (0.12) 0.04 (0.11)
-0.5 0.05 (0.15) 0.04 (0.12) 0.02 (0.12)
-0.4 0.02 (0.15) 0.03 (0.11) 0.02 (0.12)
-0.3 0.01 (0.13) 0.01 (0.11) 0.00 (0.11)
-0.2 0.00 (0.12) 0.00 (0.09) 0.00 (0.07)
-0.1 0.00 (0.10) 0.00 (0.07) 0.01 (0.06)
0.0 0.00 (0.09) 0.00 (0.05) 0.00 (0.04)
0.1 0.00 (0.10) 0.00 (0.07) 0.00 (0.06)
0.2 0.01 (0.12) 0.00 (0.09) 0.00 (0.08)
0.3 0.01 (0.13) 0.01 (0.11) 0.01 (0.10)
0.4 0.03 (0.14) 0.01 (0.12) 0.01 (0.11)
0.5 0.04 (0.16) 0.03 (0.12) 0.02 (0.11)
0.6 0.08 (0.15) 0.05 (0.12) 0.03 (0.12)
0.7 0.10 (0.14) 0.07 (0.12) 0.05 (0.11)
0.8 0.12 (0.13) 0.09 (0.10) 0.07 (0.10)
0.9 0.13 (0.11) 0.10 (0.08) 0.08 (0.09)
1.0 0.09 (0.05) 0.06 (0.03) 0.04 (0.03)
Table 6: The magnitude of empirical biases (standard errors) of $\hat{\rho}_{n,k}$ when $\alpha=5$. Optimal $k$s are selected using the method from Danielsson et al. (2016), with averages of $k=9$ ($N=100$), $k=29$ ($N=500$), and $k=79$ ($N=2000$).
$\rho_{XY}$ $N=100$ $N=500$ $N=2000$
-1.0 0.14 (0.06) 0.10 (0.04) 0.08 (0.04)
-0.9 0.20 (0.12) 0.17 (0.09) 0.13 (0.08)
-0.8 0.19 (0.14) 0.16 (0.10) 0.12 (0.10)
-0.7 0.16 (0.14) 0.13 (0.10) 0.10 (0.10)
-0.6 0.12 (0.14) 0.10 (0.11) 0.08 (0.11)
-0.5 0.09 (0.14) 0.07 (0.11) 0.06 (0.10)
-0.4 0.06 (0.14) 0.05 (0.09) 0.04 (0.09)
-0.3 0.04 (0.13) 0.03 (0.09) 0.02 (0.09)
-0.2 0.01 (0.12) 0.02 (0.08) 0.01 (0.06)
-0.1 0.00 (0.11) 0.01 (0.06) 0.00 (0.06)
0.0 0.01 (0.10) 0.00 (0.06) 0.00 (0.04)
0.1 0.01 (0.11) 0.01 (0.07) 0.01 (0.06)
0.2 0.02 (0.12) 0.01 (0.08) 0.00 (0.07)
0.3 0.03 (0.13) 0.03 (0.09) 0.02 (0.08)
0.4 0.06 (0.13) 0.05 (0.09) 0.04 (0.09)
0.5 0.09 (0.14) 0.07 (0.10) 0.06 (0.09)
0.6 0.13 (0.14) 0.10 (0.11) 0.08 (0.11)
0.7 0.16 (0.13) 0.13 (0.10) 0.10 (0.10)
0.8 0.19 (0.13) 0.15 (0.10) 0.13 (0.09)
0.9 0.21 (0.12) 0.17 (0.09) 0.14 (0.08)
1.0 0.14 (0.06) 0.11 (0.04) 0.08 (0.04)
Table 7: The magnitude of empirical biases (standard errors) of $\hat{\rho}_{n,k}$ when $\alpha=3$. Optimal $k$s are selected using the method from Clauset et al. (2009), with averages of $k=60$ ($N=100$), $k=273$ ($N=500$), and $k=991$ ($N=2000$).
$\rho_{XY}$ $N=100$ $N=500$ $N=2000$
-1.0 0.10 (0.03) 0.10 (0.02) 0.09 (0.02)
-0.9 0.15 (0.06) 0.14 (0.05) 0.13 (0.04)
-0.8 0.13 (0.08) 0.12 (0.06) 0.12 (0.05)
-0.7 0.11 (0.08) 0.10 (0.06) 0.10 (0.04)
-0.6 0.09 (0.08) 0.09 (0.06) 0.08 (0.04)
-0.5 0.07 (0.09) 0.06 (0.06) 0.06 (0.04)
-0.4 0.05 (0.08) 0.05 (0.05) 0.05 (0.04)
-0.3 0.03 (0.07) 0.03 (0.05) 0.03 (0.03)
-0.2 0.02 (0.06) 0.02 (0.03) 0.02 (0.02)
-0.1 0.01 (0.04) 0.01 (0.02) 0.01 (0.01)
0.0 0.00 (0.04) 0.00 (0.01) 0.00 (0.01)
0.1 0.01 (0.04) 0.01 (0.02) 0.01 (0.01)
0.2 0.01 (0.06) 0.02 (0.04) 0.02 (0.02)
0.3 0.03 (0.07) 0.03 (0.04) 0.03 (0.03)
0.4 0.05 (0.08) 0.05 (0.05) 0.04 (0.04)
0.5 0.07 (0.08) 0.07 (0.05) 0.06 (0.04)
0.6 0.09 (0.09) 0.09 (0.06) 0.08 (0.05)
0.7 0.11 (0.09) 0.11 (0.06) 0.10 (0.04)
0.8 0.13 (0.08) 0.13 (0.05) 0.12 (0.04)
0.9 0.14 (0.07) 0.14 (0.05) 0.13 (0.04)
1.0 0.10 (0.03) 0.10 (0.02) 0.09 (0.02)
Table 8: The magnitude of empirical biases (standard errors) of $\hat{\rho}_{n,k}$ when $\alpha=4$. Optimal $k$s are selected using the method from Clauset et al. (2009), with averages of $k=53$ ($N=100$), $k=226$ ($N=500$), and $k=791$ ($N=2000$).
$\rho_{XY}$ $N=100$ $N=500$ $N=2000$
-1.0 0.15 (0.03) 0.15 (0.02) 0.14 (0.02)
-0.9 0.22 (0.06) 0.22 (0.03) 0.21 (0.03)
-0.8 0.20 (0.06) 0.20 (0.04) 0.19 (0.03)
-0.7 0.17 (0.06) 0.17 (0.04) 0.16 (0.03)
-0.6 0.14 (0.06) 0.14 (0.04) 0.13 (0.02)
-0.5 0.11 (0.06) 0.10 (0.03) 0.10 (0.02)
-0.4 0.08 (0.06) 0.07 (0.03) 0.07 (0.02)
-0.3 0.05 (0.06) 0.05 (0.03) 0.05 (0.02)
-0.2 0.03 (0.05) 0.03 (0.02) 0.03 (0.01)
-0.1 0.01 (0.05) 0.01 (0.02) 0.01 (0.01)
0.0 0.00 (0.04) 0.00 (0.02) 0.00 (0.01)
0.1 0.01 (0.05) 0.01 (0.02) 0.01 (0.01)
0.2 0.03 (0.05) 0.03 (0.02) 0.03 (0.01)
0.3 0.05 (0.05) 0.05 (0.03) 0.05 (0.02)
0.4 0.07 (0.06) 0.08 (0.03) 0.07 (0.02)
0.5 0.11 (0.06) 0.10 (0.03) 0.10 (0.02)
0.6 0.14 (0.06) 0.14 (0.04) 0.13 (0.02)
0.7 0.17 (0.06) 0.17 (0.04) 0.16 (0.03)
0.8 0.20 (0.06) 0.20 (0.04) 0.19 (0.03)
0.9 0.22 (0.06) 0.22 (0.03) 0.21 (0.03)
1.0 0.15 (0.03) 0.15 (0.02) 0.14 (0.02)
Table 9: The magnitude of empirical biases (standard errors) of $\hat{\rho}_{n,k}$ when $\alpha=5$. Optimal $k$s are selected using the method from Clauset et al. (2009), with averages of $k=48$ ($N=100$), $k=187$ ($N=500$), and $k=602$ ($N=2000$).
$\rho_{XY}$ $N=100$ $N=500$ $N=2000$
-1.0 0.19 (0.03) 0.19 (0.02) 0.18 (0.02)
-0.9 0.28 (0.06) 0.28 (0.03) 0.27 (0.02)
-0.8 0.25 (0.06) 0.26 (0.03) 0.25 (0.02)
-0.7 0.22 (0.06) 0.21 (0.03) 0.21 (0.02)
-0.6 0.17 (0.06) 0.17 (0.03) 0.17 (0.02)
-0.5 0.14 (0.06) 0.13 (0.03) 0.13 (0.02)
-0.4 0.10 (0.06) 0.10 (0.03) 0.10 (0.02)
-0.3 0.06 (0.06) 0.07 (0.03) 0.06 (0.02)
-0.2 0.04 (0.05) 0.04 (0.03) 0.04 (0.01)
-0.1 0.02 (0.05) 0.02 (0.03) 0.02 (0.01)
0.0 0.00 (0.05) 0.00 (0.02) 0.00 (0.01)
0.1 0.01 (0.05) 0.02 (0.03) 0.02 (0.01)
0.2 0.04 (0.05) 0.04 (0.03) 0.04 (0.01)
0.3 0.07 (0.05) 0.07 (0.03) 0.06 (0.01)
0.4 0.10 (0.06) 0.10 (0.03) 0.10 (0.02)
0.5 0.14 (0.06) 0.14 (0.03) 0.13 (0.02)
0.6 0.18 (0.06) 0.17 (0.03) 0.17 (0.02)
0.7 0.22 (0.06) 0.22 (0.03) 0.21 (0.02)
0.8 0.25 (0.06) 0.25 (0.03) 0.25 (0.02)
0.9 0.28 (0.05) 0.28 (0.03) 0.27 (0.02)
1.0 0.19 (0.03) 0.19 (0.02) 0.18 (0.02)
D.2 Relation to the measures $\chi$ and $\bar{\chi}$
The joint distribution of $\|X\|$ and $\|Y\|$ can be used to assess the likelihood of extreme curves $X$ and $Y$ occurring simultaneously, where “extreme” refers to their size measured by the norm. Since $\|X\|$ and $\|Y\|$ are scalars, we can apply to them two commonly used extremal measures $\chi$ and $\bar{\chi}$ introduced by  Coles et al. (1999). This allows us to determine whether extreme $\|X\|$ and $\|Y\|$ values occur simultaneously.
We start by recalling the definitions of $\chi$ and $\bar{\chi}$. Let $F_{U}$ and $F_{V}$ are the marginal distribution functions of nonnegative random variables $U$ and $V$. The measure $\chi$ is defined as $\chi=\lim_{q\to 1}\chi(q)$, where
$\chi(q)=P(F_{U}(U)>q|F_{V}(V)>q),\quad 0<q<1.$
If $U$ and $V$ are asymptotically independent, then $\chi=0$, and if they are asymptotically dependent, then $\chi\in(0,1]$. The measure $\bar{\chi}$ is defined as $\bar{\chi}=\lim_{q\to 1}\bar{\chi}(q)$, where
$\bar{\chi}(q)=\frac{2\log P(F_{U}(U)>q)}{\log P(F_{U}(U)>q,F_{V}(V)>q)}-1,\quad 0<q<1.$
If $U$ and $V$ are asymptotically independent, then $\bar{\chi}\in[-1,1)$, and if they are asymptotically dependent, then $\bar{\chi}=1$. These two measures are thus complementary and it is useful to apply them together. We will demonstrate that our ECC is complementary to them because it provides additional information on the shapes of the extremal curves.
We generate random curves $X$ and $Y$ as described in equation (17), with $N=1000$, for $\rho_{XY}\in\{0,0.4,0.7,1\}$, and compute $\chi(q)$ and $\bar{\chi}(q)$ using $(\|X\|,\|Y\|)$. We did not consider the case when $\rho_{XY}$ is negative, as the results are similar to those for $|\rho_{XY}|$. The results are presented in Figure 7 for each value of $\rho_{XY}$. When $\rho_{XY}=0$, extreme curves in $X$ and $Y$ do not occur simultaneously, and the corresponding values for $\chi$ and $\bar{\chi}$ are both close to zero, see the upper left paired plot in Figure 7. When $\rho_{XY}\in\{0.4,0.7,1\}$, extreme curves in $X$ and $Y$ tend to occur simultaneously. In that case, $\chi$ should be greater than 0 and $\bar{\chi}(q)$ should approach 1 as $q\to 1$, which is observed in the other paired plots in Figure 7.
Figure 7: The values of $\chi(q)$ and $\bar{\chi}(q)$ (solid) with 95% confidence bands (dashed) are displayed in the paired plots. The values are computed from $\|X\|$ and $\|Y\|$, where $X$ and $Y$ are defined as in (17). The paired plots are arranged for $\rho_{XY}=0$ in the upper left, $0.4$ in the upper right, $0.7$ in the lower left, and $1$ in the lower right.
It is important to emphasize that the measures $\chi$ and $\bar{\chi}$ for $(\|X\|$, $\|Y\|$) should be used with the extremal correlation coefficient $\rho_{XY}$. Although $\chi$ and $\bar{\chi}$ quantify whether extreme curves in $X$ and $Y$ occur simultaneously, they do not account for the shapes of curves. Therefore, $\rho_{XY}$ complements $\chi$ and $\bar{\chi}$ by evaluating the relationship between the shapes of the curves, highlighting the new, functional aspect of $\rho_{XY}$. To illustrate this, consider the following toy example:
$X(t)=Z_{1}\phi_{1}(t)+N_{1}\phi_{2}(t);\ \ Y(t)=Z_{1}\phi_{2}(t)+N_{2}\phi_{1}(t),$ (33)
where $Z_{1}$, $N_{1}$, $N_{2}$, $\phi_{1}$, and $\phi_{2}$ are defined in Lemma 5.1. Since $X$ and $Y$ share $Z_{1}$, extreme events occur in both $X$ and $Y$ simultaneously. It then follows that $\chi=\bar{\chi}=1$, as shown in Fig 8. However, when examining the extreme curves in $X$ and $Y$, their patterns are unrelated since $\phi_{1}$ and $\phi_{2}$ are orthogonal. In this case, $\rho_{XY}$ captures the lack of similarity between the shapes, resulting in an estimate close to 0.
Figure 8: The values of $\chi(q)$ and $\bar{\chi}(q)$ (solid) with 95% confidence bands (dashed) are displayed in the paired plot. The values are computed from $\|X\|$ and $\|Y\|$, where $X$ and $Y$ are defined as in (33).
D.3 Effect of phase variation on $\rho_{XY}$
Phase variation occurs when some properties of curves shift over time, as seen with growth spurts occurring at different times for different individuals. Any variation in phase can affect the extremal correlation coefficient. In general, if extremal curves in the two samples are out of phase, the sample ECC will become closer to zero. For example, if heat waves tend arrive at different times at different locations, this will result in the ECC closer to zero than if the heat wave arrival times matched. To illustrate this, we consider the data generating process described in (17), but with $\phi_{k}$ in $Y(t)$, replaced by
$\phi_{k}^{*}(t)=0,\ {\rm if}\ t\leq 0.3,\quad\phi_{k}^{*}(t)=\phi_{k}\left(t-0.3\right),\ {\rm if}\ t>0.3.$
Table 10 presents the $\hat{\rho}_{n,k}$ with and without phase shift. The results indicate that phase shift brings the $\hat{\rho}_{n,k}$ closer to zero.
Table 10: The $\hat{\rho}_{n,k}$ for samples without and with phase variation when $\alpha=3$ and $N=100$. The cut-off $k$s is selected using the method of Danielsson et al. (2016).
Phase variation
$\rho_{XY}$ No Yes
-1.0 -0.96 -0.90
-0.9 -0.84 -0.78
-0.8 -0.73 -0.69
-0.7 -0.66 -0.62
-0.6 -0.57 -0.54
-0.5 -0.49 -0.47
-0.4 -0.41 -0.39
-0.3 -0.31 -0.29
-0.2 -0.22 -0.20
-0.1 -0.11 -0.09
0.0 0.00 -0.02
0.1 0.11 0.10
0.2 0.21 0.20
0.3 0.32 0.28
0.4 0.40 0.38
0.5 0.50 0.46
0.6 0.56 0.53
0.7 0.66 0.61
0.8 0.74 0.69
0.9 0.84 0.78
1.0 0.96 0.90
[…]
Appendix E Supplementary results for Section 6
This section presents Hill plots for the CIDRs ETF data discussed in Section 6.1 and the temperature data discussed in Section 6.2. These plots suggest that the extreme curves discussed in those sections appear to be regularly varying, as the plots generally show stable regions.
Figure 9: Hill plots of the norm of centered CIDRs for each sector ETF, with the Hill estimates (solid) and 95% confidence intervals (dashed). From left to right, the upper row shows: XLY, XLP, XLE; the middle row: XLF, XLV, XLI; and the lower row: XLB, XLK, XLU.
Figure 10: Hill plots of the norm of centered temperature curves for each location, with the Hill estimates (solid) and 95% confidence intervals (dashed). From left to right, Fort Collins, CO, Colorado Springs, CO, and Austin, TX.